A Bibliography of ACM SIGMICRO Newsletter

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 April 2021 Version 1.00

Title word cross-reference

#3 [37]. #4 [38].

$17.95 [110]. 2n [200]. 3 [42]. $52.95 [487].

-D [42]. -way [200].

0 [135]. 0-89006-064-9 [135]. 000 [354]. 0471-50336-3 [487].

1 [148, 112, 114, 234, 240]. 11/60 [295]. 11/780 [336, 286]. 16-bit [207]. 17 [71]. 1700 [113, 161]. 19 [375]. 1971 [27]. 1976 [99]. 1980 [215, 214, 247]. 1981 [244, 243, 248]. 1982 [245]. 1983 [311]. 1988 [445]. 19th [376].

21 [413, 412, 415, 414]. 2100 [169]. 2100A [97]. 22 [479, 416]. 23 [482]. 27 [310]. 286 [309]. 2nd [100].

3000 [104]. 32-000 [354]. 32-bit [372]. 3220 [226]. 360 [140]. 370 [136].

400 [181]. 4341 [213].

60 [295].

72 [97]. 764 [28]. 780 [336, 286]. 794 [26].

8000 [202, 203, 198, 199]. 8600 [388].

9 [135].

Abstract [459, 335]. abstraction [209]. Abstracts [58, 63, 69, 86, 488]. Academic [99, 317]. acceleration [358]. access [516, 312]. ACM [99, 467, 84]. across [130]. acyclic [507]. adaptable [476]. Adaptation [225]. additional [2]. address [182, 517, 279, 300, 144, 250]. addressed [516]. advanced [290]. Advances [352, 311]. ago [10]. Agrawala [99]. aided [405, 422, 192]. Algebra [161]. Algorithm [194, 328, 424, 130, 23, 332, 374, 378, 451, 144]. Algorithmic [387]. Algorithms [123, 486]. Alice [257]. allocation [144]. allowing [47]. Alternative [337, 111]. alternatives [109]. ALU [508]. ALUs [248]. AMD2900 [224]. AN/UYK [71]. AN/UYK-17 [71]. analysis [292, 399, 114, 229, 400, 154, 487]. analyzer [11]. Andrews [214]. annotated [45]. annual [78, 376]. antibiotic [405]. APL [235]. Application [180, 30, 453, 91, 245]. Applications [319, 83, 491, 99]. Applying [308]. approach [422, 120, 285, 6, 512, 377, 344, 418, 351, 124]. Architectural [25, 185]. Architecture [99, 304, 318, 336, 202, 271, 452, 500, 94, 251, 111, 252, 116, 225, 420, 432, 476, 262, 423, 150, 239, 293, 504, 190, 191, 349, 499, 355, 462]. Architectures [446, 147, 347, 353, 461, 454, 338]. area [326]. arithmetic [286, 467, 170, 71, 166]. array [153]. arrays [455]. Art [487]. Artech [135, 311]. articles [61, 68, 72, 218, 242]. ’AS888 [372]. ’AS888/’AS890 [372]. ’AS890 [372]. Ashok [99]. ASIC [478]. aspects [100]. assembled [83]. assembler [212, 368, 290, 228, 224, 104]. assignment [147, 300]. assist [136]. assistant [403]. assisted [471]. Association [73]. associative [292, 89]. attributed [336]. automata [9]. automated [302, 6, 197]. Automatic [109, 157, 237, 300, 345, 334]. automation [411, 7, 422]. awards [479]. Axiomatic [296].

B [27]. B1700 [101, 151]. B1726 [113]. B1800 [120, 313]. based [396, 329, 500, 148, 399, 225, 209, 398, 463, 400, 430, 485, 428, 334, 466, 382]. basic [151]. BBMSS [398]. behavioral [424]. better [280, 24]. between [98, 373, 88]. Beux [246]. bibliography [45]. big [207]. binding [200]. bipolar [83]. BIT [215, 401, 448, 207, 350, 372, 398, 289, 108, 247, 248]. BIT-Slice [215, 401, 350, 289, 247, 248]. bit-splice [398]. block [204]. blocks [17, 83]. Board [480]. Book [487, 445, 247]. books [95]. boundaries [130]. Branch [494, 504]. branch-intensive [504]. branches [263, 510, 518]. branching [505]. Brick [247]. Bridging [98]. bring [24]. buffers [516]. building [17, 204]. built [285]. Burroughs [79, 120, 113, 151]. Bylaws [481].

C [26, 28, 483, 184]. C-20 [26, 28]. C-string [184]. caches [516]. Call [482, 370]. candidates [332]. capability [396, 382]. capability-based [396, 382]. Carlo [266]. case [292, 353, 130, 354, 340, 439, 392]. caused [377]. certain [109]. chair [415, 414]. Characterisation [241]. Characteristics [27, 251, 491]. checking [457, 456]. chip [320, 138]. chips [224, 50]. choices [474]. Choust [246]. Chroust [216]. Cincinnati [274]. circuit [83]. CISCs [287]. Classes [131]. classical [272]. clocked [473]. Code [514, 509, 147, 114, 224, 266, 469, 491]. coded [161]. codes [277]. Coding [14]. collapsing [508]. Colorade [249]. Colorado [367]. column [10, 12, 15]. combination [429]. Combining [454]. commenting [119]. communication [390]. communications [316]. compacted [302]. Compacting [426, 511]. Compaction [327, 328, 486, 424, 283, 282, 303, 281, 469, 377, 265, 485, 230, 374, 378, 240, 325, 364, 425, 272, 468]. compaction-based [485]. Company [247, 216]. comparison [147, 142, 373]. compatible [402]. compilation [437, 454, 428]. compiler [483, 77, 184, 471, 495, 427, 496, 177, 343, 485, 312, 264, 359]. compiler-assisted [471]. compilers [450, 409]. Compiling [354]. complement [168]. components [206]. comprehensive [506]. compressed [499]. Compressing [205]. Computer [245, 405, 487, 214, 423, 85, 47, 217, 396, 314, 284, 327, 147, 132, 422, 94, 188, 347, 353, 123, 41, 317, 115, 67, 6, 192, 42, 33, 225, 121, 76, 35, 34, 465, 410, 149, 201, 190, 191, 218, 29, 513]. computer-aided [192]. Computers [27, 28, 24, 472, 243]. Computing [38, 37, 185, 239, 459]. concept [42]. Conception [202]. Concepts [38, 8]. concurrency [441, 429]. concurrent [288, 503, 449]. concurrently [269, 337]. conditional [437, 505, 510]. conflicts [88]. conjugate [492]. Considerations [240, 148, 65, 190, 191]. considered [90]. considering [124]. constant [264]. constrained [497]. constraints [425]. constructs [315, 174]. Contemporaey [38]. contemporary [8]. context [281]. Contrasting [222]. Control [260, 243, 28, 324, 103, 285, 114, 205, 82, 32, 369, 88, 458, 439, 269, 136, 91, 36, 122, 366, 349, 174, 214]. controlled [241]. controller [103, 316]. Controllers [248]. Controlling [502]. controls [7, 67]. convenience [24]. conventional [410]. converter [342]. copyright [402]. core [476]. Cornell [371]. corporate [79]. Corporation [79]. correct [431]. correctness [341]. Cost [453, 390, 518]. Cost-effective [453]. coupling [460, 407, 264]. CP [306]. CPU [372]. Critical [357]. cross [104]. crossbar [316]. Crossing [268]. cultures [258]. current [301, 13]. curriculum [33, 34]. customized [366]. CVAX [440].

D [243, 42]. dangerous [90]. Data [390, 503, 143, 358, 500, 103, 82, 209, 377]. database [121]. death [401]. debugger [397]. debugging [350, 323]. December [38]. decision [443]. decisions [339, 267, 213]. Dedham [135]. Defensive [259]. defies [287]. Definition [467, 64, 265]. delayed [263]. delays [505]. density [114]. dependence [498]. dependent [232]. derivation [196]. descent [336]. description [180, 253, 446, 229, 50, 385, 231]. descriptions [424]. Design [148, 347, 339, 103, 395, 215, 32, 204, 65, 164, 210, 178, 269, 456, 128, 322, 451, 243, 149, 248, 107, 383, 411, 70, 422, 302, 285, 317, 224, 192, 372, 261, 109, 453, 475, 428, 230, 152, 307, 278, 190, 191, 366, 359, 93, 167, 247]. designers [24]. destructible [241]. details [279]. detecting [88]. deterministic [495]. developing [422, 299]. Development [227, 314, 384, 254, 329, 402, 408, 198, 111, 321, 80, 361, 183, 456, 289, 382, 189, 256, 310, 362, 447]. device [186]. devices [402]. DGCAS [189]. diagnostic [350]. diagnostic/debugging [350]. differences [434]. Digital [67, 172, 11, 215]. dimension [435]. Direct [185, 184, 94, 28, 452, 188]. Direct-execution [94]. directed [501, 255, 461]. display [42, 257]. Distributed [419]. Division [417]. DOAS [462]. documentation [278]. Donnamarie [248]. during [2]. Dynamic [118, 507, 489, 335, 42, 47, 511, 397, 294]. dynamic-programming [511]. dynamically [115, 201].

E-Machine [291]. early [45]. easily [239]. EASY [112]. ECLIPSE [132, 188, 105, 106]. ECLIPSE(R) [202, 203, 198, 199]. economics [206]. Editor [445, 2, 210, 12, 15]. educational [406]. effect [402]. Effective [496, 494, 200, 453]. Efficient [348, 386, 500, 497, 176, 378, 144]. effort [363]. Efrem [135, 311]. electronic [413]. element [71]. elementary [467]. elements [117]. eliminate [469]. Elmer [226]. embedded [499]. EMMY [140]. EMMY/360 [140]. employs [307]. empty [10]. emulates [306]. Emulating [271].

Emulation [276, 28, 172, 148, 89, 186, 75, 169, 125, 8, 140, 38]. emulations [97]. emulator [111, 149, 318]. Emulators [379, 131]. engine [289]. engineer [96]. Engineering [214, 445, 249, 370, 222, 367, 274, 323, 447]. Enhanced [498, 510]. enhancement [92, 126]. enumeration [109]. environment [384, 118, 400, 465]. environments [172]. Epsilon [483]. equivalence [59]. error [457]. Euclid [449]. Euromicro [73, 86, 219]. European [73]. evaluating [409]. evaluation [41, 477, 109, 466, 270, 351, 201]. events [430]. eventually [24]. example [469]. Executing [499]. execution [147, 452, 184, 94, 188, 496, 315, 504, 196, 506]. exercise [317]. exhaustive [328]. Experience [160, 299, 151, 312, 179, 152]. experiment [187]. experimental [75, 312]. Experimentation [284]. experiments [364]. expert [298]. exploitation [494]. Exploiting [492, 441, 512]. expression [23, 31]. Extended [457, 136]. extensible [115]. extension [391]. extraction [429].

F [38]. faces [24]. facility [181, 288, 313, 136]. Facility/370 [136]. fast [418]. faster [211]. Fault [16]. feature [14]. fetch [471, 59, 506]. fetches [348]. file [492]. files [514]. finite [9]. Firmware [159, 214, 249, 445, 367, 418, 274, 126, 130, 192, 267, 40, 222, 385, 262, 334, 319, 152, 256, 216, 246]. ‘FLEX’ [223]. flexible [178, 476]. floating [286, 502]. floating-point [502]. Flow [297, 358, 390, 344]. flows [241]. foresighted [468]. formal [345, 39, 377, 431]. Formalization [197]. Formally [399]. formation [328]. formats [124]. FORTRAN [147]. Forward [471]. found [257]. Foundations [99]. framework [486, 453]. FTSC [158]. Functional [450, 20, 473, 27]. functionally [402]. Functions [246, 312]. fundamentals [192]. future [301].

G [99, 216, 246]. gap [98]. Garland [248]. General [169, 236, 32]. generalization [281]. generalized [101]. generating [177, 175]. generation [407, 345, 77, 329, 156, 233, 344, 389, 509, 237, 264]. generator [386, 236, 128]. genetic [486]. George [243]. Glenford [215]. Global [344, 436, 174, 283, 281, 497, 193, 230, 374, 325, 364, 145]. Graph [428, 498, 390, 297, 344]. graphics [42, 29]. graphs [241, 507]. Group [84]. Growth [37]. GURPR [436].

H [27]. Habib [445]. halfbaked [48]. Hall [245, 243]. hard [108]. Hardware [358, 434, 183, 221, 64, 120, 180, 517, 200, 348, 213, 141, 139, 96, 136, 152, 323, 330, 429, 355, 216]. hardware/microcode [213, 330]. hardware/software [120]. Harry [110]. HDLs [422]. heavily [471, 502]. Heuristics [193]. Hewlett [169]. Hierarchical [175, 142]. hierarchies [139]. High [211, 301, 235, 232, 320, 143, 299, 360, 234, 164, 233, 343, 357, 239, 179, 208, 397, 117]. High-level [235, 232, 360, 233, 343, 397]. high-performance [320]. higher [46]. higher-level [46]. highly [212, 503, 491]. highly-parallel [212]. highly-vectorized [491]. HILEVEL [408]. Hill [110, 247]. hint [516]. HLL [333, 315, 102, 152]. HLLDA [287]. HLLDAs [287]. Holland [244, 216, 246]. horizontal [384, 407, 441, 469, 374, 239, 300, 144]. horizontally [237]. host [251, 150, 338]. House [135, 311]. HP [97]. HPS [357, 356, 389, 442]. HPSm [441]. hybrid [120, 387]. hyperblock [496]. hypothetical [410].

I/O [187, 186]. I370 [435]. iAPX [309]. IBM [213, 136]. IC [50]. IDAS [411]. idea [48]. Identification [293]. IEEE [26, 27, 28, 286, 370]. II [31, 50, 45]. Image [338]. impact [345, 381]. Implementation [143, 185, 336, 202, 320, 194, 115, 388, 333, 346, 116, 387, 178, 474, 378, 199, 138, 442, 355]. implementations [223]. implemented [151]. Implementing [432, 337, 246]. implications [337, 20]. Improved [328, 315, 256]. improvement [85, 325, 201]. Improving [139, 295]. increase [281]. increased [508]. inc(R) [67]. Incremental [468]. independent [360, 176, 380]. index [172]. industry [54]. influencing [339]. Information [461, 404, 406, 281]. inherent [458]. innovations [218]. instantiation [234]. Instruction [59, 129, 513, 440, 182, 328, 516, 471, 492, 494, 200, 280, 51, 455, 512, 508, 238, 379, 389, 472, 124, 491, 518, 250, 146, 506]. instruction-level [492, 494, 512, 508, 491]. instructional [466]. instructions [292]. instruments [379]. Integers [417]. integrated [411, 83, 6]. integrating [333]. Intel [104]. intellectual [363]. intensive [504]. Interactive [181, 350, 42, 151, 178, 359]. Interdata [62]. Interest [37, 84]. interface [57, 268]. interfaces [515]. Interlisp [160]. Interlock [508]. intermediate [188]. interpretation [409, 102]. interpreter [101, 224, 449, 418, 330, 313]. interpretive [139]. interrupt [132]. Introduction [28, 162, 139, 356]. investigation [489]. IRELAND [416]. ISA [409, 318]. ISBN [487, 135]. isolation [16]. Issue [488]. Issues [230, 357, 503].

J [215, 216]. Jain [487]. JAM [365]. James [247]. John [487, 247]. journal [61, 68, 72]. journals [56, 87]. Jr [110]. JRS [411]. Judgement [444]. July [26, 27, 28]. jump [200]. jumps [437, 348]. just [365].

Katzan [110]. Keynote [279, 250]. keyword [121]. Kleir [26]. Kraft [243].

L [246]. laboratory [40, 466, 93, 276, 186]. landmark [443]. language [107, 211, 232, 163, 162, 255, 299, 23, 50, 346, 116, 313, 461, 234, 209, 66, 233, 343, 385, 5, 176, 423, 230, 410, 231, 330, 296, 362, 117]. language-directed [255]. languages [450, 46, 188, 180, 65, 164, 208, 462]. large [25]. LASS [141]. law [402]. Lawson [27]. legal [443]. Lessor [28]. level [284, 327, 211, 232, 46, 492, 494, 301, 114, 299, 235, 360, 234, 512, 164, 508, 177, 233, 343, 439, 230, 226, 179, 208, 201, 491, 397, 167, 117]. Levels [150]. Limitation [493]. limitations [180]. limited [514]. limits [490]. link [103, 210]. LISP [113, 418]. literature [13]. Local [283, 328, 139, 240, 468]. logic [358, 289, 431, 215]. long [10]. lookahead [485]. loop [500, 464, 463]. loops [437, 509, 293, 504, 511, 510]. Louisiana [277]. low [230, 201, 397]. low-level [201, 397]. LSI [215, 137, 306]. LSI-CP [306].

M [37, 224]. M29 [290]. Machine [136, 232, 339, 395, 251, 113, 151, 137, 223, 457, 369, 387, 297, 268, 176, 410, 451, 380, 338, 503, 296, 291]. machine-dependent [232]. machine-independent [176]. machine-specific [296]. machines [150, 459]. Macintosh [465]. mailbag [413]. mainframe [432, 307]. maintenance [6]. makes [62]. Mallach [135, 311]. managing [103]. manual [36]. many [24]. Mapping [315]. MASCO [317]. Mass [135]. mathematical [431]. MATHILDA [81]. matrices [91]. maximize [309]. may [24]. MBALM [113]. MBALM/1700 [113]. mcASM [373, 368]. McGraw [110, 247]. McGraw-Hill [110, 247]. MCU [36]. MDS [256]. MDSL [107]. measure [238]. measurement [120, 85]. measurements [451]. mechanism [506]. medical [406]. medium [177, 19]. meeting [480, 57]. meetings [57]. memory [101, 515, 89, 470, 20, 106]. merging [494]. meta [212, 224]. meta-assembler [212, 224]. META29R [373]. metasimulator [331]. METASTEP [362]. Method [180, 471, 200, 340, 436]. methodology [153, 456]. methods [344, 431, 429, 445]. Metrics [129]. Michael [214]. Mick [247]. MICOS [142]. Micro [166, 155, 202, 161, 80, 183, 91, 127, 146, 375, 482, 413, 412, 479, 415, 416, 414]. MICRO-19 [375]. Micro-21 [413, 412, 415, 414]. Micro-22 [479, 416]. Micro-23 [482]. micro-architecture [202]. micro-coded [161]. micro-instruction [146]. micro-operations [146]. micro-programmed [183]. micro-programming [91, 127]. Micro8 [228]. Microarchitecture [507, 253, 474, 396, 440, 286, 320, 358, 339, 477, 255, 475, 385, 357, 356, 309, 442]. microarchitectures [473, 409, 377, 337]. microassembler [119, 105]. Microassembly [326]. Microbe [119]. microblock [265]. Microcode [403, 361, 340, 265, 363, 382, 425, 217, 384, 450, 407, 336, 328, 386, 486, 254, 424, 426, 345, 324, 329, 368, 290, 408, 354, 198, 399, 388, 228, 299, 448, 346, 283, 427, 443, 213, 457, 303, 409, 377, 341, 177, 233, 400, 430, 210, 315, 297, 344, 428, 326, 334, 230, 378, 108, 431, 237, 323, 351, 330, 380, 325, 364, 352, 196, 264, 272]. microcode-based [400]. microcoded [182, 394, 331, 316, 137, 156]. Microcomputer [321, 107, 324, 322, 28]. microcomputers [83]. microcontrol [221]. microcontrollers [157]. MicrodataNews [74]. microdiagnostics [456, 199]. microgrammed [19]. microinstruction [200, 125, 231]. microinstructions [241, 389]. micromachine [179]. microoperations [345, 88]. Microprocessing [242, 100, 387, 244, 73]. microprocessing/microprogramming [387]. Microprocessor [307, 247, 133, 134, 440, 207, 252, 441, 183, 456, 138, 493, 306]. Microprocessors [246, 260, 478, 205, 308]. Microprogram [214, 369, 278, 167, 117, 46, 181, 229, 398, 236, 178, 128, 270, 189, 195, 342, 175, 124, 18, 158, 9]. Microprogram/ [9]. Microprogrammable [133, 134, 190, 191, 212, 143, 17, 115, 204, 293, 149, 201, 349]. Microprogrammed [292, 383, 129, 446, 154, 36, 138, 122, 11, 16, 284, 132, 142, 422, 347, 353, 120, 402, 160, 103, 194, 285, 123, 41, 113, 205, 23, 31, 83, 42, 32, 116, 361, 121, 449, 261, 458, 269, 289, 19, 237, 366, 106, 243]. Microprogrammed-system [383]. Microprogramming [4, 52, 70, 99, 38, 33, 21, 90, 455, 35, 96, 216, 371, 71, 43, 242, 275, 29, 305, 78, 419, 421, 232, 165, 1, 92, 335, 64, 100, 30, 98, 163, 301, 162, 180, 288, 317, 227, 255, 118, 24, 333, 235, 360, 22, 249, 45, 51, 75, 234, 387, 209, 40, 95, 65, 66, 435, 164, 233, 343, 125, 392, 367, 5, 262, 274, 466, 176, 102, 230, 8, 34, 60, 226, 465, 410, 295, 298, 208, 85, 47, 393, 258, 381, 20, 376, 93, 296, 259, 244, 135, 37, 84, 110, 311, 445]. microprograms [327, 77, 302, 39, 282, 197, 193, 374, 300, 144, 145, 359, 397, 174, 26]. Microsequencer [203, 262, 365]. MICROSIM [125]. microstore [211]. MicroTAL [232]. MICROTEC [368]. MICSIM [23, 31]. MIDAS [229]. MIDDLE [162]. MIDETT [383]. MIDL [231]. MIES [475]. Migration [333, 292, 335, 130, 267, 332, 334, 294]. military [306]. MIMD [271]. MIMOLA [426, 428]. mind [278]. minicomputer [169]. minicomputers [16, 97]. Minutes [480]. MIPS [252]. MISC [513]. MIXER [298]. MLM [395]. model [473, 477, 303, 88]. models [297]. Modifications [286]. Modifying [517]. Modula [313]. modular [262, 334, 174, 217]. module [155]. modulo [509, 510]. monitor [267]. monitoring [313]. Monograph [99]. Monte [266]. Motivation [486]. much [24].

Mühlbacher [216]. multifunction [438]. Multilingual [27, 395]. multimicroprocessor [148]. Multiple [470, 14, 318, 513]. multiplication [168]. multiprocessor [284, 327, 17, 316, 109]. multithreaded [512]. multiway [348]. MV [202, 203, 198, 199]. MV/8000 [202, 203, 198, 199]. Myers [215].

N [243]. nanocode [240]. nanodata [234, 240, 114]. NCR [354]. NCR/32 [354]. NCR/32-000 [354]. needs [24]. network [316]. News [49, 82, 44, 219]. newsletters [86]. No [26, 28]. Noguez [246]. non [495]. non-deterministic [495]. noninvasive [400]. Norman [135, 311]. North [244, 216, 246]. North-Holland [244, 216, 246]. Nostrand [445]. notation [10, 46]. Note [417, 170, 173, 108]. Notes [54, 56, 55, 57, 53, 87, 2]. notices [49, 44]. novel [91]. NP [108]. NP-hard [108]. nucleus [142].

O [187, 186]. object [462]. objectives [164]. occupancy [124]. October [100]. off [213]. ongoing [446]. Operating [92, 126, 142, 112, 267, 430]. Operation [438, 470]. operational [251]. operations [467, 166, 146]. operator [502]. operators [283]. optimal [464, 504]. optimization [460, 427, 266, 463, 193, 108, 195, 124, 145, 174, 26]. optimized [77]. Optimizing [263, 264, 359]. oracle [238]. organization [47]. oriented [185, 335, 162, 293, 149, 294, 462]. Other [22, 220, 409]. Our [137]. overview [81]. own [62].

P [246]. PACE [270]. package [101]. Packard [169]. packing [146]. pages [487]. PAISLey [446]. papers [484, 482, 100, 58, 370, 220]. parable [10]. Parallel [378, 28, 212, 451]. parallelism [492, 494, 512, 508, 238, 490, 201]. parallelization [464]. parallelizing [485]. parametric [89]. parsing [440, 23, 31]. part [48, 23, 31, 50]. partitioning [390]. PASC [152]. PASC-HLL [152]. Pascal [295]. Patchable [324]. path [503]. PDP [295]. PDP-11 [295]. PDP-11/60 [295]. pedagogical [35, 43]. Peephole [460]. Performance [487, 201, 172, 320, 142, 489, 313, 139, 400, 357, 239, 295, 451, 309, 307, 85, 493]. Perkin [226]. Perkin-Elmer [226]. personalization [225]. perspective [272]. pescription [345]. Phase [407, 264]. photos [375]. physically [516]. physically-addressed [516]. pictorial [412]. pin [517]. Pipe [505]. pipeline [153, 505]. pipelined [471, 433, 457, 502, 472]. pipelines [438]. pipelining [495, 437, 392, 319, 308, 391, 436]. PLA [326]. PLAs [221]. point [286, 502]. Positive [417]. potential [238]. power [62]. pp [38, 37, 26, 28]. practical [351]. practice [180]. practices [98, 21, 22]. pre [348]. pre-fetches [348]. predicated [496]. prefetching [501]. Preliminary [404]. preloading [500]. Prentice [245, 243]. Prentice-Hall [245]. Press [99, 214, 248]. Price [487]. PRIM [111]. Prime [181]. Primer [110]. primitives [48]. principles [98, 21, 22, 214]. Probability [245]. problem [14, 293, 272]. problems [377]. procedure [103]. processing [155, 433, 392, 71, 122]. processor [153, 515, 42, 137, 156, 141, 432, 19, 382, 250, 167, 506, 27]. processor/memory [515]. processors [52, 350, 143, 471, 422, 515, 434, 331, 501, 32, 361, 514, 497, 453, 19, 237, 306]. product [83]. Program [91, 498, 59, 41, 80, 415, 136, 490]. programmable [206, 114]. programmed [183]. Programming [156, 153, 333, 379, 91, 36, 166, 511, 127]. programs [384, 211, 406, 238, 150, 499]. project [446, 277]. projects [371]. Prolog [452, 339, 354, 388, 369, 337, 338]. proof [296]. Proposal [144]. proposals [337]. prospective [379]. prospects [301]. protection [363, 218]. prototype [323]. pseudofields [469]. public [79]. Publication [3]. Publishing [216]. PUMPKIN [66]. purpose [32, 169].

QM [234, 148, 112, 114, 234, 240]. QM-1 [234, 148, 112, 114, 234, 240]. quality [304]. quasi [349]. Queueing [245].

R [216, 28]. Raj [487]. rational [170, 166]. rationale [356]. Rauscher [99]. realization [395]. reconfigurable [438]. recursive [336]. reduce [516, 517, 390, 161]. REDUCE/1700 [161]. reduced [324]. reduction [283, 236, 326]. reference [36]. Reflective [97]. regarding [357, 337]. region [498]. register [492, 48, 514, 423]. registers [518]. Reinhold [445]. related [218]. relations [79]. relevant [100]. Reliability [245]. Reliable [243]. Remamoorthy [26]. reordering [472]. Report [165, 376]. representation [150]. request [303]. required [147]. requirements [185, 517]. requirements-oriented [185]. Research [108, 367, 371]. resembling [46]. resolves [14]. resource [303, 497, 124]. resource-constrained [497]. resources [377]. respective [337]. Response [171, 173]. Restructurable [216]. results [292]. retargetable [386, 290, 448, 427, 233, 343, 344, 428]. retargeting [385]. Review [245, 244, 99, 135, 38, 37, 26, 215, 214, 445, 216, 246, 27, 8, 243, 110, 247, 248, 311, 28, 3, 487, 100, 57, 22]. reviewers [55]. Reviews [13]. Richter [246]. RISC [394, 287, 499]. RISCs [287]. risk [324]. Robert [38]. role [198, 51, 34]. ROM [205]. Rosin [38]. routine [132]. rules [296]. Run [389, 211, 420]. Run-time [389, 420]. runs [41].

S [245]. S* [353, 234]. S*-system [353]. S.J.C.C. [1]. sample [182]. Sampling [484]. scalar [501, 491]. SCAMP [404]. SCARCE [453]. schedulable [239]. scheduled [509]. scheduler [495]. scheduling [498, 507, 489, 427, 514, 281, 463, 497, 239, 325, 438, 510]. schema [163, 509]. scheme [492, 234, 505, 387, 334]. schemes [260, 221]. Science [245, 214, 406, 33, 35, 34]. scientific [239, 491]. SDLC [103]. SDVS [340, 341]. secure [462]. selection [332]. self [285, 123, 119, 456]. self-checking [456]. self-tuning [123]. semantic [471]. seminar [75]. September [37]. sequencing [51, 334, 502]. Sequential [338]. Series [99]. service [132]. Session [279, 2, 1]. set [320, 48, 250]. Sets [129, 143, 280, 236]. seventh [78]. sharing [118, 226]. shuffling [469]. SIGMICRO [480, 57, 481]. signal [52, 155, 392, 71, 122]. significance [60]. SIGPLAN [57]. SIGPLAN/SIGMICRO [57]. simple [76]. simulate [410]. simulation [107, 358, 41, 346, 116, 398, 423]. simulator [329, 23, 31, 125, 178, 179, 18]. simulator-based [329]. single [238, 138]. six [147]. sized [19]. Slice [215, 350, 401, 448, 289, 247, 248]. SLIM [346]. Small [243]. Smith [27]. smoothability [490]. sociology [393]. Soft [314]. Software [99, 321, 436, 182, 159, 130, 495, 120, 437, 198, 448, 80, 370, 222, 392, 183, 391, 217, 352, 429]. software-firmware [130]. Some [2, 239, 195, 364]. Son [215]. Sondak [135, 311]. Sons [487]. source [313]. Southwestern [277]. space [147]. SPAM [430]. Special [84]. specific [194, 453, 296]. specification [347]. Specified [446]. spectrum [154]. speculative [506]. speech [156]. speed [143, 315]. speeds [314]. splice [398]. SRDAG [281]. Stanford [186, 276, 140]. Stanley [445]. start [273]. started [24]. State [249, 367, 387]. statements [147, 46]. static [399]. station [310]. Statistical [41]. Statistics [245]. Step [447, 362, 310]. Step-27 [310]. store [324, 366]. stores [439]. STPM [248]. Strategies [26]. strategy [102]. Stream [513, 238, 389]. streams [472]. Stride [501]. string [184]. Structure [17, 28, 185, 194, 477]. Structured [368, 70, 51, 334, 349]. structures [470, 461, 117]. studies [195]. study [292, 353, 130, 354, 340, 439, 392, 19]. studying [305]. submitted [484]. subsystem [350]. suite [409]. summary [413, 412]. superscalar [515, 514, 497, 493]. supplies [62]. Support [80, 286, 507, 335, 267, 496, 136, 262, 466, 289, 217, 352]. supporting [152, 506, 462]. Survey [406, 133, 134, 208]. Surveys [38, 37]. symbolic [196]. symposium [100]. syntax [235]. synthesis [197, 222, 297, 157, 293]. System [215, 314, 383, 411, 421, 254, 92, 126, 7, 142, 422, 206, 148, 353, 120, 160, 408, 112, 115, 111, 161, 229, 267, 81, 261, 398, 341, 233, 430, 428, 466, 289, 298, 270, 256, 175, 362, 447, 217, 140]. System/360 [140]. Systems [487, 172, 25, 212, 329, 24, 379, 351]. systolic [455].

table [463]. tailored [24]. taken [2]. Target [360]. Target-independent [360]. targeting [460]. task [507]. TDL [330]. Teaching [40, 125, 35, 465]. technical [413]. technique [383, 437, 14, 463, 497, 392, 454, 145, 511]. techniques [489, 59, 253, 266, 326, 239, 152, 308, 307]. technology [30]. Ten [127, 211]. terminal [257]. test [285, 330]. testable [269]. testbed [181]. tester [323]. texts [184]. their [337]. Theoretical [100]. theory [39]. therapy [405]. there [257]. thoughts [287]. three [207]. time [147, 516, 118, 420, 389, 349]. time-transparent [349]. times [211]. timing [425]. today [207]. Tomlinson [99]. tool [460, 305, 182, 430, 35, 475, 189, 43]. tools [383, 198, 227, 321, 448, 183, 385, 382, 362]. total [256]. Toy [243]. Trace [427, 281, 325]. tracing [430]. trade [213]. trade-off [213]. Tradeoffs [515]. Transactions [26, 27, 28, 370]. transfer [48, 423]. transformation [121]. transient [377]. transition [91]. Translation [516, 222]. Transparent [335, 349]. Tree [282]. trend [90]. trends [25]. Trivedi [245]. Tuning [420, 123, 442]. Two [223, 284, 327, 114, 439]. two-level [284, 327, 439]. Two’s [168]. Two’s-complement [168]. type [335, 294]. type-oriented [294].

UCSD [295]. UDSYS [254].

unaugmented [458]. unification [451, 355]. unit [285, 433, 121, 71, 36, 355]. units [269, 366]. universal [150, 342]. universities [53]. University [249, 277, 367, 371, 274, 275]. UNIX [267, 226]. untestability [458]. URCR [391]. URPR [392, 391]. usage [301]. use [111, 224, 281]. used [321]. user [115, 227, 448, 24, 204, 226, 47]. user-extensible [115]. user-microprogrammable [204]. Using [485, 238, 431, 486, 132, 358, 467, 17, 59, 354, 317, 299, 446, 151, 283, 340, 169, 469, 496, 236, 439, 392, 453, 221, 423, 465, 518]. UYK-17 [71].

V [37, 71, 28]. V1 [38, 37]. V73 [128]. validation [181]. Vanderbilt [275]. VARIAN [128, 97]. various [489]. VAX [336, 286, 320, 324, 388, 321, 389, 474, 322, 442]. VAX-11 [336, 286]. VAX-11/780 [336, 286]. vectorized [491]. Venice [100].

Verification [353, 158, 159, 162, 340, 222, 322]. verifying [341]. Version [368, 469]. versus [159]. vertical [292, 335, 267, 441, 294]. very [397]. VHDL [424]. via [388, 441, 265, 91, 295]. video [257]. view [261]. Virtual [187, 136, 101, 106]. visual [421]. VLIW [514, 497, 453, 454, 504]. VLIW/superscalar [514]. VLSI [260, 324, 321, 205, 346, 225, 261, 476, 466, 322, 381, 318, 306]. VLSI-based [225]. VM [517, 136]. VM/370 [136]. VMX [251]. vocoder [11]. Vol [26, 28].

W [27]. waveform [156]. way [200]. Where [273]. White [248]. width [236]. Wiley [487, 215]. Wilkes [37]. window [485]. Wing [243]. without [505]. word [88]. words [146]. workbench [291]. Workshop [4, 78, 165, 376]. write [431]. writers [101].

XB [71]. XB-1 [71].

Y-Pipe [505]. years [127]. York [487, 110].

zero [518]. zero-cost [518].

References

Black:1969:MSJ

[1] Colin Black. The microprogramming session at the S.J.C.C. ACM SIGMI-CRO Newsletter, 1(2):3–4, August 1969. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/2579798.2579799.

Anonymous:1969:SAN

[2] Anonymous. Some additional notes taken by your editor during the session. ACM SIGMICRO Newsletter, 1(2):5, August 1969. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579798.2579800.

Anonymous:1969:PR

[3] Anonymous. Publication review. ACM SIGMICRO Newsletter, 1(2):6, August 1969. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/2579798.
2579801.

Anonymous:1969:WM

[4] Anonymous. Workshop on microprogramming. ACM SIGMICRO Newsletter, 1(2):7–35, August 1969. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579798.2579802.

Osborne:1969:LM

[5] Thomas E. Osborne. A language for microprogramming. ACM SIGMICRO Newsletter, 1(2):36–48, August 1969. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/2579798.2579803.

Hackl:1970:IAA

[6] F. J. Hackl and R. W. Shirk. An integrated approach to automated computer maintenance. ACM SIGMICRO Newsletter, 1(3):7, January 1970. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579804.2579805.

Buckingham:1970:CAS

[7] B. R. S. Buckingham, W. C. Carter, W. R. Crawford, and G. A. Nowell. The controls automation system. ACM SIGMICRO Newsletter, 1(3):8, January 1970. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579804.2579806.

Rosin:1970:RCC

[8] Robert F. Rosin. Review: contemporary concepts of microprogramming and emulation. ACM SIGMICRO Newsletter, 1(3):9, January 1970. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579804.2579807.

Matsevityy:1970:MFA

[9] L. V. Matsevityy. Microprogram/ finite automata. ACM SIGMICRO Newsletter, 1(3):10, January 1970. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579804.2579808.

Anonymous:1970:ECP

[10] Anonymous. The empty column: a parable about a “new notation” of long ago. ACM SIGMICRO Newsletter, 1(4):5–6, March 1970. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579809.2579810.

Ancona:1970:MDV

[11] Enrico I. Ancona. A microprogrammed digital vocoder analyzer. ACM SIG-MICRO Newsletter, 1(4):10–29, March 1970. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/2579809.
2579811.

Anonymous:1971:EC

[12] Anonymous. Editor’s column. ACM SIGMICRO Newsletter, 2(1):15–16, April 1971. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/2579812.
2579813.

Husson:1971:RCL

[13] Samir S. Husson. Reviews of current literature. ACM SIGMICRO Newsletter, 2(1):17–18, April 1971. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579812.2579814.

Galey:1971:CTR

[14] J. Michael Galey. Coding technique resolves multiple feature problem. ACM SIGMICRO Newsletter, 2(1):19–23, April 1971. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2579812.2579815.

Merwin:1971:EC

[15] R. E. Merwin. Editor’s column. ACM SIGMICRO Newsletter, 2(3):1–2, October 1971. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/2580713.
2580714.

Anonymous:1971:FIM

[16] Anonymous. Fault isolation in microprogrammed minicomputers. ACM SIG-MICRO Newsletter, 2(3):3–26, October 1971. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/2580713.
2580715.

Davis:1971:SMU

[17] R. L. Davis and S. Zucker. Structure of a multiprocessor using microprogrammable building blocks. ACM SIGMICRO Newsletter, 2(3):27–42, October 1971. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/2580713.
2580716.

Young:1971:MS

[18] Steve Young. A microprogram simulator. ACM SIGMICRO Newsletter, 2(3):43–56, October 1971. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2580713.2580717.

Redfield:1971:SMP

[19] Stephen R. Redfield. A study in microprogrammed processors: a medium sized microgrammed processor. ACM SIGMICRO Newsletter, 2(3):57–58, October 1971. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/2580713.
2580718.

Tucker:1971:FMM

[20] S. G. Tucker. Functional memory and its microprogramming implications. ACM SIGMICRO Newsletter, 2(3):58–59, October 1971. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2580713.2580719.

Husson:1971:MPP

[21] Samir S. Husson and Paula Schneider. Microprogramming: principles and practices. ACM SIGMICRO Newsletter, 2(3):59–60, October 1971. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2580713.2580720.

Husson:1971:ORM

[22] Samir S. Husson. Other review of microprogramming: principles and practices. ACM SIGMICRO Newsletter, 2(3):61–63, October 1971. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/2580713.2580721.

Habib:1972:MMEa

[23] S. Habib. MICSIM — a microprogrammed expression parsing simulator: Part I— language and algorithm. ACM SIGMICRO Newsletter, 2(4):7–11, January 1972. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1316527.1316528.

Haavind:1972:MFM

[24] Robert C. Haavind. The many faces of microprogramming: what started out as a convenience for systems designers may eventually bring computers much better tailored to users’ needs. ACM SIGMICRO Newsletter, 2(4):12–16, January 1972. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1316527.1316529.

Amdahl:1972:ATL

[25] Gene M. Amdahl. Architectural trends in large systems. ACM SIGMICRO Newsletter, 2(4):17–29, January 1972. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1316527.1316530.

Hellerman:1972:ROS

[26] Leo Hellerman. Review of “Optimization Strategies for Microprograms, by Kleir and Remamoorthy”, IEEE Transactions, Vol. C-20 No. 7, July 1971, pp. 783–794. ACM SIGMICRO Newsletter, 2(4):30–31, January 1972. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1316527.1316531.

Roberts:1972:RFC

[27] John D. Roberts. Review of “Functional Characteristics of a Multilingual Processor by H. W. Lawson and B. K. Smith”, IEEE Transactions on Computers, July 1971. ACM SIGMICRO Newsletter, 2(4):31–33, January 1972. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1316527.1316532.

Tsuchiya:1972:RID

[28] Masahiro Tsuchiya. Review of “An Introduction to the Direct Emulation of Control Structure by a Parallel Microcomputer, by V. R. Lessor”, IEEE Transactions on Computers, Vol. C-20 No. 7, July 1971, pp. 751– 764. ACM SIGMICRO Newsletter, 2(4):33–35, January 1972. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1316527.1316533.

vanDam:1972:MCG

[29] Andries van Dam. Microprogramming for computer graphics. ACM SIGMI-CRO Newsletter, 3(1):3–7, April 1972. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1316535.1316536.

Clapp:1972:AMT

[30] J. A. Clapp. The application of microprogramming technology. ACM SIGMI-CRO Newsletter, 3(1):8–47, April 1972. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1316535.1316537.

Habib:1972:MMEb

[31] S. Habib. MICSIM — a microprogrammed expression parsing simulator: Part II — the simulator. ACM SIGMI-CRO Newsletter, 3(1):48–56, April 1972. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1316535.1316538.

Hoff:1972:DMC

[32] George Hoff. Design of microprogrammed control for general purpose processors. ACM SIGMICRO Newsletter, 3(2):57–64, July 1972. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1096507.1096508.

Hemming:1972:MCS

[33] C. W. Hemming. Microprogramming in the computer science curriculum. ACM SIGMICRO Newsletter, 3(3):6, October 1972. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1316540.1316541.

Rosin:1972:RMC

[34] R. F. Rosin and G. Frieder. The role of microprogramming in the computer science curriculum. ACM SIGMICRO Newsletter, 3(3):7–10, October 1972. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1316540.1316542.

Merwin:1972:MPT

[35] R. E. Merwin. Microprogramming as a pedagogical tool in teaching computer science. ACM SIGMICRO Newsletter, 3(3):11–17, October 1972. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1316540.1316543.

Roberts:1972:MCU

[36] John D. Roberts, J. Ihnat, and W. R. Smith. Microprogrammed control unit (MCU) programming reference manual. ACM SIGMICRO Newsletter, 3 (3):18–57, October 1972. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1316540.1316544.

Gorman:1972:RGI

[37] T. Gorman. Review of “The Growth of Interest in Microprogramming by M. V. Wilkes,” Computing Surveys, V1, #3, pp. 139–145 (September 1969). ACM SIGMICRO Newsletter, 3(3):58, October 1972. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1316540.
1316545.

Gorman:1972:RCC

[38] T. Gorman. Review of “Contemporaey Concepts of Microprogramming and Emulation by Robert F. Rosin,” Computing Surveys, V1, #4, pp. 197–212 (December 1969). ACM SIGMICRO Newsletter, 3(3):58, October 1972. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1316540.1316546.

Ito:1973:TFM

[39] T. Ito. A theory of formal microprograms. ACM SIGMICRO Newsletter, 4(1):5–17, April 1973. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217116.1217117.

Lewis:1973:TMF

[40] T. G. Lewis, A. R. DeKock, and H. R. Alcorn. Teaching microprogramming: a firmware laboratory. ACM SIGMICRO Newsletter, 4(1):18–23, April 1973. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217116.1217118.

Farber:1973:SEP

[41] Hans-Jürgen Färber. Statistical evaluation of program runs on a microprogrammed computer through simulation. ACM SIGMICRO Newsletter, 4(1):24–29, April 1973. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217116.1217119.

Hartenstein:1973:MDP

[42] Reiner W. Hartenstein and Klaus D. Mueller. A microprogrammed display processor concept for 3-D dynamic interactive computer graphics. ACM SIGMI-CRO Newsletter, 4(1):30–35, April 1973. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217116.1217120.

Tassinari:1973:MPT

[43] Alberto F. Tassinari. Microprogramming: a pedagogical tool. ACM SIGMI-CRO Newsletter, 4(1):36–38, April 1973. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217116.1217121.

Merwin:1973:NN

[44] R. E. Merwin. News and notices. ACM SIGMICRO Newsletter, 4(1):39– 50, April 1973. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217116.
1217122.

Jones:1973:ABM

[45] Louise H. Jones and Kenneth Carvin. An annotated bibliography on microprogramming II: early 1972 — early 1973. ACM SIGMICRO Newsletter, 4(2):7–18, July 1973. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217124.1217127.

Berndt:1973:MNR

[46] von H. Berndt. A microprogram notation resembling statements of higherlevel languages. ACM SIGMICRO Newsletter, 4(2):19–27, July 1973. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217124.1217128.

Thomas:1973:COA

[47] Richard T. Thomas. Computer organization for allowing dynamic user microprogramming. ACM SIGMICRO Newsletter, 4(2):28–42, July 1973. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217124.1217129.

Hartenstein:1973:HIA

[48] Reiner W. Hartenstein. A halfbaked idea about a set of register transfer primitives: (part 1). ACM SIGMICRO Newsletter, 4(2):43–52, July 1973. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217124.1217130.

Habib:1973:NN

[49] Stanley Habib. News and notices. ACM SIGMICRO Newsletter, 4(2):53– 63, July 1973. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1217124.
                                        1217125.

Hartenstein:1973:TLD

[50] Reiner W. Hartenstein. Towards a language for the description of IC chips: (Part II). ACM SIGMICRO Newsletter, 4(3):6–16, October 1973. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217132.1217133.

Jones:1973:RIS

[51] Louise H. Jones. The role of instruction sequencing in structured microprogramming. ACM SIGMICRO Newsletter, 4(3):17–21, October 1973. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217132.1217134.

Ashcraft:1973:MSP

[52] W. D. Ashcraft. Microprogramming of signal processors. ACM SIGMICRO Newsletter, 4(3):22–25, October 1973. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217132.1217135.

Habib:1973:NU

[53] Stanley Habib. Notes from universities. ACM SIGMICRO Newsletter, 4(3):26–28, October 1973. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217132.1217136.

Habib:1973:NI

[54] Stanley Habib. Notes from industry. ACM SIGMICRO Newsletter, 4(3):29, October 1973. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217132.
1217137.

Habib:1973:NR

[55] Stanley Habib. Notes from reviewers. ACM SIGMICRO Newsletter, 4 (3):30–37, October 1973. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217132.1217138.

Habib:1973:NJ

[56] Stanley Habib. Notes from journals. ACM SIGMICRO Newsletter, 4(3):38–45, October 1973. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217132.1217139.

Habib:1973:NRM

[57] Stanley Habib. Notes from recent meetings: a review of the SIGPLAN/ SIGMICRO interface meeting. ACM SIGMICRO Newsletter, 4(3):46–58, October 1973. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217132.
1217140.

Habib:1974:ARP

[58] Stanley Habib. Abstracts from recent papers. ACM SIGMICRO Newsletter, 4(4):4–5, January 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217142.1217143.

Eastwood:1974:IFT

[59] Douglas E. Eastwood. Instruction fetch techniques using program equivalence. ACM SIGMICRO Newsletter, 4(4):6–23, January 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217142.1217144.

Rosin:1974:SM

[60] Robert F. Rosin. The significance of microprogramming. ACM SIGMICRO Newsletter, 4(4):24–39, January 1974. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217142.1217145.

Habib:1974:RJAa

[61] Stanley Habib. Recent journal articles. ACM SIGMICRO Newsletter, 4(4):40–48, January 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217142.1217146.

Habib:1974:INM

[62] Stanley Habib. Interdata now makes own power supplies. ACM SIGMICRO Newsletter, 4(4):49–60, January 1974. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217142.1217147.

Habib:1974:RAa

[63] Stanley Habib. Recent abstracts. ACM SIGMICRO Newsletter, 5(1):3–10, April 1974. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1217149.
                                        1217150.

Chroust:1974:HDM

[64] Gerhard Chroust. A hardware definition of microprogramming. ACM SIGMI-CRO Newsletter, 5(1):11–14, April 1974. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217149.1217151.

Lloyd:1974:DCM

[65] Gregory R. Lloyd and Andries van Dam. Design considerations for microprogramming languages. ACM SIGMICRO Newsletter, 5(1):15–44, April 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217149.1217152.

Lloyd:1974:PAM

[66] Gregory R. Lloyd. PUMPKIN: (another) microprogramming language. ACM SIGMICRO Newsletter, 5(1):45– 76, April 1974. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217149.
1217153.

Habib:1974:DCC

[67] Stanley Habib. Digital computer controls inc(R). ACM SIGMICRO Newsletter, 5(1):77–83, April 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217149.1217154.

Habib:1974:RJAb

[68] Stanley Habib. Recent journal articles. ACM SIGMICRO Newsletter, 5(1):84–87, April 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217149.1217155.

Habib:1974:RAb

[69] Stanley Habib. Recent abstracts. ACM SIGMICRO Newsletter, 5(2):5–6, April 1974. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217157.
1217158.

Boulaye:1974:MSD

[70] Guy Boulaye. Microprogramming and structured design. ACM SIGMICRO Newsletter, 5(2):19–28, April 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217157.1217159.

Rauscher:1974:MUX

[71] Tomlinson G. Rauscher. Microprogramming the AN/UYK-17(XB-1)(V) signal processing element signal processing arithmetic unit. ACM SIGMICRO Newsletter, 5(2):29–63, April 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217157.1217160.

Habib:1974:RJAc

[72] Stanley Habib. Recent journal articles. ACM SIGMICRO Newsletter, 5(2):64–66, April 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217157.1217161.

Habib:1974:EEA

[73] Stanley Habib. Euromicro: the European Association for Microprocessing. ACM SIGMICRO Newsletter, 5(2):67–69, April 1974. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217157.1217162.

Habib:1974:M

[74] Stanley Habib. MicrodataNews. ACM SIGMICRO Newsletter, 5(2):70–78, April 1974. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217157.
1217163.

Jones:1974:ESM

[75] Louise H. Jones. An experimental seminar on microprogramming and emulation. ACM SIGMICRO Newsletter, 5(2):79–83, April 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217157.1217164.

McKeeman:1974:SC

[76] W. M. McKeeman. A simple computer. ACM SIGMICRO Newsletter, 5(3):16–48, October 1974. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096510.1096511.

Blain:1974:CGO

[77] G. Blain, M. Perrone, and N. X. Hong. A compiler for the generation of optimized microprograms. ACM SIGMICRO Newsletter, 5(3):49–67, October 1974. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096510.1096512.

Anonymous:1974:SAW

[78] Anonymous. The seventh annual workshop on microprogramming. ACM SIG-MICRO Newsletter, 5(3):68–89, October 1974. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096510.
1096513.

Anonymous:1974:BCC

[79] Anonymous. Burroughs Corporation: corporate public relations. ACM SIG-MICRO Newsletter, 5(3):93–144, October 1974. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096510.
1096514.

Hodges:1975:SSM

[80] Bobby C. Hodges and Andrew J. Edwards. Support software for micro program development. ACM SIGMICRO Newsletter, 5(4):17–24, January 1975. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217166.1217167.

Kornerup:1975:OMS

[81] Peter Kornerup and Bruce D. Shriver. An overview of the MATHILDA system. ACM SIGMICRO Newsletter, 5(4):25–53, January 1975. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217166.1217168.

Habib:1975:NCD

[82] Stanley Habib. News control data. ACM SIGMICRO Newsletter, 5(4):54–61, January 1975. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1217166.
                                        1217169.

Habib:1975:NPA

[83] Stanley Habib. New product applications: microprogrammed bipolar microcomputers are assembled from integrated circuit blocks. ACM SIGMICRO Newsletter, 5(4):62–70, January 1975. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217166.1217170.

Habib:1975:ASI

[84] Stanley Habib. ACM Special Interest Group on Microprogramming. ACM SIGMICRO Newsletter, 6(1):5–6, April 1975. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217172.
1217173.

Snyder:1975:CPI

[85] David C. Snyder. Computer performance improvement by measurement and microprogramming. ACM SIGMI-CRO Newsletter, 6(1):7–14, April 1975. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217172.1217174.

Habib:1975:RAE

[86] Stanley Habib. Recent abstracts from Euromicro newsletters. ACM SIGMI-CRO Newsletter, 6(1):15–17, April 1975. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217172.1217175.

Habib:1975:NJ

[87] Stanley Habib. Notes from journals. ACM SIGMICRO Newsletter, 6(1):24–25, April 1975. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217172.1217176.

Lewis:1975:CWM

[88] Ted Lewis and Bruce Shriver. A control word model for detecting conflicts between microoperations. ACM SIG-MICRO Newsletter, 6(3):6–9, September 1975. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217180.
1217181.

Dalrymple:1975:PAM

[89] S. H. Dalrymple. A parametric associative memory emulation. ACM SIG-MICRO Newsletter, 6(3):10–36, September 1975. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217180.
1217182.

Lehman:1975:MTC

[90] M. M. Lehman. Microprogramming trend considered dangerous. ACM SIG-MICRO Newsletter, 6(3):37–39, September 1975. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217180.
1217183.

Oliver:1976:PCT

[91] A. Ron Oliver and Neil D. Jones. Program control via transition matrices: a novel application of microprogramming. ACM SIGMICRO Newsletter, 7(1):8–27, March 1976. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217202.1217203.

Brown:1976:OSE

[92] George E. Brown, Richard H. Eckhouse, and Robert P. Goldberg. Operating system enhancement through microprogramming. ACM SIGMICRO Newsletter, 7(1):28–33, March 1976. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217202.1217204.

Vickery:1976:MDL

[93] Christopher Vickery. A microprogramming design laboratory. ACM SIG-MICRO Newsletter, 7(1):34–48, March 1976. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217202.
1217205.

Chu:1976:DEC

[94] Yaohan Chu. Direct-execution computer architecture. ACM SIGMICRO Newsletter, 7(1):49–53, March 1976. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217202.1217206.

Lewis:1976:RBM

[95] T. G. Lewis and Bruce Shriver. Recent books on microprogramming. ACM SIGMICRO Newsletter, 7(1):54, March 1976. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217202.
1217207.

Mick:1976:MHE

[96] John R. Mick. Microprogramming for the hardware engineer. ACM SIGMI-CRO Newsletter, 7(2):17–25, June 1976. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217209.1217210.

Ehlers:1976:REH

[97] Bryan L. Ehlers, Gary L. Harmon, and Larry L. Wear. Reflective emulations of the HP 2100A and Varian 72 minicomputers. ACM SIGMICRO Newsletter, 7(2):26–42, June 1976. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217209.1217211.

Cohen:1976:BGB

[98] David Cohen and Ming T. Liu. Bridging the gap between principles and practices in microprogramming. ACM SIGMI-CRO Newsletter, 7(2):43–60, June 1976. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217209.1217212.

Chattergy:1976:RFM

[99] Rahul Chattergy. Review of “Foundations of Microprogramming; Architecture, Software, and Applications by Ashok K. Agrawala and Tomlinson G. Rauscher,” Academic Press, ACM Monograph Series, 1976. ACM SIGMI-CRO Newsletter, 7(2):61–63, June 1976. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217209.1217214.

Chroust:1976:TAM

[100] G. Chroust. Theoretical aspects of microprogramming: a review of relevant papers at the 2nd symposium on microprocessing and microprogramming (Venice, October 1976). ACM SIG-MICRO Newsletter, 7(4):29–30, December 1976. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1217218.
1217219.

Belgard:1976:GVM

[101] Richard A. Belgard. A generalized virtual memory package for B1700 interpreter writers. ACM SIGMICRO Newsletter, 7(4):31–45, December 1976. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217218.1217220.

Petit:1976:MSH

[102] J. Petit, D. Litaize, B. Lecussan, and J. P. Sansonnet. A microprogramming strategy for HLL interpretation. ACM SIGMICRO Newsletter, 7(4):46–69, December 1976. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217218.1217221.

Dromard:1976:DMC

[103] D. Dromard and D. Lafage. Design of a microprogrammed controller managing the SDLC data link control procedure. ACM SIGMICRO Newsletter, 7(4):70–86, December 1976. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217218.1217222.

Willen:1976:ICA

[104] David Willen. An Intel 3000 cross assembler. ACM SIGMICRO Newsletter, 7(4):87–94, December 1976. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217218.1217223.

Martin:1977:EM

[105] Daniel Martin. An Eclipse microassembler. ACM SIGMICRO Newsletter, 8(1):13–23, March 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217225.1217226.

dHautcourt-Carette:1977:MVM

[106] Françoise d’Hautcourt Carette. A microprogrammed virtual memory for Eclipse. ACM SIGMICRO Newsletter, 8(2):10–20, June 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217228.1217229.

Adamowicz:1977:MMD

[107] Michael Adamowicz and Jamshed Mirza. MDSL: a microcomputer design and simulation language. ACM SIGMI-CRO Newsletter, 8(2):21–39, June 1977. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217228.1217230.

Robertson:1977:RNM

[108] Edward L. Robertson. Research note: microcode bit optimization is NPhard. ACM SIGMICRO Newsletter, 8(2):40–43, June 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217228.1217231.

Middleton:1977:AEE

[109] Tony Middleton. Automatic enumeration and evaluation of certain multiprocessor design alternatives. ACM SIGMI-CRO Newsletter, 8(2):44–48, June 1977. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1217228.1217232.

Thomas:1977:RMP

[110] Richard T. Thomas. Review of “Microprogramming Primer by Harry Katzan Jr.,” McGraw-Hill, New York, $17.95. ACM SIGMICRO Newsletter, 8(2):49–51, June 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1217228.1217234.

Goldberg:1977:PSA

[111] Joel Goldberg, Alvin Cooperband, and Louis Gallenson. The PRIM system: an alternative architecture for emulator development and use. ACM SIGMICRO Newsletter, 8(3):1–6, September 1977. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1014196.803309.

Flink:1977:EOS

[112] Charles W. Flink. EASY — an operating system for the QM-1. ACM SIG-MICRO Newsletter, 8(3):7–14, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014196.
803310.

Griss:1977:MML

[113] M. L. Griss and M. R. Swanson.

MBALM/1700: a microprogrammed LISP machine for the Burroughs B1726. ACM SIGMICRO Newsletter, 8(3): 15–25, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014196.803311.

Frieder:1977:ACD

[114] Gideon Frieder and Jill Miller. An analysis of code density for the two level programmable control of the Nanodata QM-1. ACM SIGMICRO Newsletter, 8(3):26–32, September 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014196.803312.

Fung:1977:IUE

[115] Fergus K. Fung and Willis K. King.

The implementation of a user-extensible system on a dynamically microprogrammable computer. ACM SIGMI-CRO Newsletter, 8(3):33–36, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014196.
803313.

Hopkins:1977:MIA

[116] William C. Hopkins and Gary Davidian. A microprogrammed implementation of an architecture simulation language. ACM SIGMICRO Newsletter, 8(3):37–46, September 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014196.803314.

Zimmermann:1977:MSH

[117] Gerhard Zimmermann. Microprogram structures for high level language elements. ACM SIGMICRO Newsletter, 8(3):47–54, September 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014196.803315.

Guha:1977:DMT

[118] Ratan K. Guha. Dynamic microprogramming in a time sharing environment. ACM SIGMICRO Newsletter, 8(3):55–60, September 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014196.803316.

Laws:1977:MSC

[119] Ben A. Laws. Microbe: a self commenting microassembler. ACM SIGMICRO Newsletter, 8(3):61–65, September 1977. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1014196.803317.

Denny:1977:BBM

[120] W. Michael Denny. The Burroughs B1800 microprogrammed measurement system: a hybrid hardware/software approach. ACM SIGMICRO Newsletter, 8(3):66–70, September 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014196.803318.

Kannan:1977:MKT

[121] Krishnamurthi Kannan, David K. Hsiao, and Douglas S. Kerr. A microprogrammed keyword transformation unit for a database computer. ACM SIG-MICRO Newsletter, 8(3):71–79, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014196.
803319.

Swartzlander:1977:MCS

[122] Earl E. Swartzlander. Microprogrammed control for signal processing. ACM SIGMICRO Newsletter, 8(3):80–84, September 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014196.803320.

El-Ayat:1977:AST

[123] K. A. El-Ayat and J. A. Howard. Algorithms for a self-tuning microprogrammed computer. ACM SIGMICRO Newsletter, 8(3):85–91, September 1977. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1014196.803321.

Tokoro:1977:AMO

[124] Mario Tokoro, Eiji Tamura, Kazuhiko Takase, and Kiichiro Tamaru. An approach to microprogram optimization considering resource occupancy and instruction formats. ACM SIGMI-CRO Newsletter, 8(3):92–108, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014196.
803322.

Mathur:1977:MMS

[125] Francis P. Mathur. MICROSIM: a microinstruction simulator for teaching microprogramming and emulation. ACM SIGMICRO Newsletter, 8(3): 109–118, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014196.803323.

Brown:1977:OSE

[126] George E. Brown, Richard Eckhouse, and Jay Estabrook. Operating system enhancement through firmware. ACM SIGMICRO Newsletter, 8(3): 119–133, September 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014196.803324.

Wilkes:1977:TYM

[127] M. V. Wilkes. Ten years and more of micro-programming. ACM SIGMICRO Newsletter, 8(4):11–13, December 1977. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1096518.1096519.

Persson:1977:DMG

[128] Magnus Persson. Design of a microprogram generator for the VAR-IAN V73. ACM SIGMICRO Newsletter, 8(4):14–20, December 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1096518.1096520.

Frisina:1977:MMI

[129] J. N. Frisina. Metrics for microprogrammed instruction sets. ACM SIG-MICRO Newsletter, 8(4):21–23, December 1977. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1096518.
                                        1096521.

Davidson:1977:CSM

[130] Scott Davidson. A case study of the migration of an algorithm across softwarefirmware boundaries. ACM SIGMICRO Newsletter, 8(4):24–33, December 1977. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096518.1096522.

Flynn:1977:CE

[131] Michael J. Flynn. Classes of emulators. ACM SIGMICRO Newsletter, 8(4):34–35, December 1977. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096518.1096523.

Brown:1978:MIS

[132] Gerry A. Brown. A microprogrammed interrupt service routine using the ECLIPSE computer. ACM SIGMICRO Newsletter, 9(1):18–22, March 1978. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096525.1096526.

Adams:1978:MMSa

[133] Phillip M. Adams. Microprogrammable microprocessor survey. ACM SIGMI-CRO Newsletter, 9(1):23–25, March 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096525.
1096527.

Adams:1978:MMSb

[134] Phillip M. Adams. Microprogrammable microprocessor survey. ACM SIGMI-CRO Newsletter, 9(2):7–38, June 1978. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096529.1096530.

Eckhouse:1978:RMN

[135] R. E. Eckhouse. Review of “Microprogramming, by Norman Sondak and Efrem Mallach”, Artech House, Dedham, Mass., 1977, ISBN 0-89006-064-9. ACM SIGMICRO Newsletter, 9(3):5, September 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096532.1096533.

Olbert:1978:ECP

[136] Arthur G. Olbert. Extended control program support: VM/370: a hardware assist for the IBM Virtual Machine Facility/370. ACM SIGMICRO Newsletter, 9(3):8–25, September 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096532.1096534.

Johannsen:1978:OMM

[137] Dave Johannsen. Our machine, a microcoded LSI processor. ACM SIG-MICRO Newsletter, 9(4):1–7, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014198.
804298.

Stritter:1978:MIS

[138] Skip Stritter and Nick Tredennick. Microprogrammed implementation of a single chip microprocessor. ACM SIG-MICRO Newsletter, 9(4):8–16, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014198.
804299.

Manthey:1978:IPI

[139] Michael J. Manthey. Improving the performance of interpretive hierarchies by the introduction of local hardware. ACM SIGMICRO Newsletter, 9 (4):17–18, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804300.

Wallach:1978:EES

[140] Walter A. Wallach. EMMY/360: an emulation of System/360 for the Stanford EMMY. ACM SIGMICRO Newsletter, 9(4):19–24, December 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804301.

Kunz:1978:LHP

[141] Paul F. Kunz, Richard N. Fall, Michael F., and Hanoch Brafman. The LASS hardware processor. ACM SIGMICRO Newsletter, 9(4):25–32, December 1978. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1014198.804302.

Burkhardt:1978:MMH

[142] Walter H. Burkhardt and Helmut E.

Maier. MICOS: a microprogrammed hierarchical operating system nucleus and its performance comparison. ACM SIGMICRO Newsletter, 9(4):33, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014198.
804303.

Caplin:1978:IHS

[143] Glenn N. Caplin, Anne R. Clayton, and Richard L. Stuart. Implementation of high speed data sets with microprogrammable data processors. ACM SIGMICRO Newsletter, 9 (4):34–39, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804304.

Tanaka:1978:PEA

[144] Tadashi Tanaka, Tadamichi Kawada, and Teiji Emori. Proposal on efficient address allocation algorithm for horizontal microprograms. ACM SIGMICRO Newsletter, 9(4):40, December 1978. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1014198.804305.

Tokoro:1978:TGO

[145] Mario Tokoro, Takashi Takizuka, Eiji Tamura, and Ichiro Yamaura. A technique of global optimization of microprograms. ACM SIGMICRO Newsletter, 9(4):41–50, December 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804306.

Wood:1978:PMO

[146] Graham Wood. On the packing of micro-operations into micro-instruction words. ACM SIGMICRO Newsletter, 9(4):51–55, December 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804307.

Belgard:1978:CCS

[147] Richard A. Belgard and Victor B. Schneider. A comparison of the code space and execution time required for FORTRAN assignment statements on six computer architectures. ACM SIGMICRO Newsletter, 9(4): 56–64, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804308.

Crocker:1978:DCQ

[148] Steve Crocker. Design considerations for a QM-1 based multimicroprocessor emulation system. ACM SIGMICRO Newsletter, 9(4):65, December 1978.

CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1014198.804309.

Soh:1978:DEO

[149] Jin W. Soh and Ron Marko. Design of an emulator oriented microprogrammable computer. ACM SIGMICRO Newsletter, 9(4):66, December 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804310.

Rau:1978:LRP

[150] B. Ramakrishna Rau. Levels of representation of programs and the architecture of universal host machines. ACM SIGMICRO Newsletter, 9(4): 67–79, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804311.

Hemphill:1978:EIB

[151] John M. Hemphill and Richard T.

Thomas. Experience with an interactive basic machine implemented using the Burroughs B1700. ACM SIGMICRO Newsletter, 9(4):80, December 1978. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1014198.804312.

Schoellkopf:1978:PHE

[152] J. P. Schoellkopf. PASC-HLL: an experience in design techniques for firmware and the supporting hardware. ACM SIGMICRO Newsletter, 9(4):81, December 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804313.

Cohen:1978:MPP

[153] Danny Cohen. A methodology for programming a pipeline array processor. ACM SIGMICRO Newsletter, 9 (4):82–89, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014198.804314.

Mulrooney:1978:MSA

[154] Timothy J. Mulrooney. Microprogrammed spectrum analysis. ACM SIG-MICRO Newsletter, 9(4):90–99, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014198.
                                        804315.

Bell:1978:MSP

[155] David M. Bell and Larry E. Hand. A micro signal processing module. ACM SIGMICRO Newsletter, 9(4):100, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014198.
                                        804316.

Kopec:1978:PMP

[156] Gary E. Kopec and Glen S. Miranker. Programming a microcoded processor for speech waveform generation. ACM SIGMICRO Newsletter, 9(4):101–111, December 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014198.804317.

Nagle:1978:ASM

[157] Andrew W. Nagle. Automatic synthesis of microcontrollers. ACM SIGMI-CRO Newsletter, 9(4):112–117, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014198.
804318.

van-Mierop:1978:VFM

[158] Dono van Mierop, Leo Marcus, and Steve Crocker. Verification of the FTSC microprogram. ACM SIGMICRO Newsletter, 9(4):118, December 1978. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1014198.804319.

Budkowski:1978:FVS

[159] S. Budkowski and P. Dembinski.

Firmware versus software verification. ACM SIGMICRO Newsletter, 9(4): 119–127, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804320.

Deutsch:1978:EMI

[160] L. Peter Deutsch. Experience with a microprogrammed Interlisp system. ACM SIGMICRO Newsletter, 9(4): 128–129, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804321.

Griss:1978:RMC

[161] Martin L. Griss and Robert R. Kessler.

REDUCE/1700: a micro-coded algebra system. ACM SIGMICRO Newsletter, 9(4):130–138, December 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804322.

Dembinski:1978:IVO

[162] P. Dembinski and S. Budkowski. An introduction to the verification oriented microprogramming language “MID-DLE”. ACM SIGMICRO Newsletter, 9(4):139–143, December 1978. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014198.804323.

Dasgupta:1978:TML

[163] Subrata Dasgupta. Towards a microprogramming language schema. ACM SIG-MICRO Newsletter, 9(4):144–153, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014198.
804324.

Malik:1978:DOH

[164] Kamran Malik and Ted Lewis. Design objectives for high level microprogramming languages. ACM SIGMI-CRO Newsletter, 9(4):154–160, December 1978. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014198.
804325.

Berg:1979:RMW

[165] Helmut K. Berg and Udo Schloms. Report on a microprogramming workshop. ACM SIGMICRO Newsletter, 10(1):6–9, March 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217236.1217237.

Thacker:1979:MPR

[166] William I. Thacker and G. W. Gorsline. Micro programming rational arithmetic operations. ACM SIGMICRO Newsletter, 10(1):10–13, March 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217236.1217238. See critical

comments [170] and response [171].

Wood:1979:MDP

[167] Graham Wood. Microprogram design at the processor level. ACM SIGMICRO Newsletter, 10(1):14–20, March 1979. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217236.1217239.

Sheue:1979:TCM

[168] A. E. Sheue. Two’s-complement multiplication. ACM SIGMICRO Newsletter, 10(1):21–23, March 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1217236.1217240.

Lewis:1979:GPE

[169] Daniel W. Lewis. General purpose emulation using the Hewlett–Packard 2100 minicomputer. ACM SIGMICRO Newsletter, 10(1):24–31, March 1979. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1217236.1217241.

Kornerup:1979:NRA

[170] Peter Kornerup. A note on rational arithmetic. ACM SIGMICRO Newsletter, 10(2):28, June 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218003.1218004. See [166,

171].

Thacker:1979:R

[171] W. I. Thacker and G. W. Gorsline.

Response. ACM SIGMICRO Newsletter, 10(2):29, June 1979. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218003.1218005. See [166,

Akonteh:1979:PIE

[172] Ayola N. Akonteh. A performance index for emulation environments in digital systems. ACM SIGMICRO Newsletter, 10(3):6–13, September 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1218007.1218008.

Kornerup:1979:NRN

[173] Peter Kornerup. A note to a response to a note. ACM SIGMICRO Newsletter, 10(3):14–15, September 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1218007.1218009.

Wood:1979:GOM

[174] Graham Wood. Global optimization of microprograms through modular control constructs. ACM SIGMICRO Newsletter, 10(4):1–6, December 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014188.803006.

Tamura:1979:HMG

[175] Eiji Tamura and Mario Tokoro. Hierarchical microprogram generating system. ACM SIGMICRO Newsletter, 10(4):7–21, December 1979. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014188.803007.

Patterson:1979:TEM

[176] David A. Patterson, Karl Lew, and Richard Tuck. Towards an efficient, machine-independent language for microprogramming. ACM SIGMI-CRO Newsletter, 10(4):22–35, December 1979. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014188.
803008.

Marti:1979:MLC

[177] Jed B. Marti and Robert R. Kessler. A medium level compiler generating microcode. ACM SIGMICRO Newsletter, 10(4):36–41, December 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014188.803009.

Mezzalama:1979:DIF

[178] M. Mezzalama and P. Prinetto. Design and implementation of a flexible and interactive microprogram simulator. ACM SIGMICRO Newsletter, 10(4):42–48, December 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014188.803010.

Schleimer:1979:EHL

[179] S. Schleimer and W. J. Meyers. Experience with a high level micromachine simulator. ACM SIGMICRO Newsletter, 10(4):49–54, December 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014188.803011.

Drongowski:1979:AHD

[180] Paul J. Drongowski and Charles W.

Rose. Application of hardware description languages to microprogramming: Method, practice, and limitations. ACM SIGMICRO Newsletter, 10(4):55–60, December 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014188.803012.

Crews:1979:IMV

[181] Phillip Crews and Laura Marie Leventhal. Interactive microprogram validation: a Prime 400 testbed facility. ACM SIGMICRO Newsletter, 10 (4):61–67, December 1979. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014188.803013.

Armbruster:1979:MTS

[182] C. Edward Armbruster. A microcoded tool to sample the software instruction address. ACM SIGMICRO Newsletter, 10(4):68–72, December 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014188.803014.

Nash:1979:HST

[183] James Nash and Mike Spak. Hardware and software tools for the development of a micro-programmed microprocessor. ACM SIGMICRO Newsletter, 10(4):73–83, December 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014188.803015.

Bridges:1979:DEC

[184] Charles W. Bridges and Abd-Elfattah Mohamed Abd-alla. Direct execution of C-string compiler texts. ACM SIGMI-CRO Newsletter, 10(4):84–92, December 1979. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014188.
                                        803016.

Anastas:1979:DAI

[185] M. S. Anastas and R. F. Vaughan. Direct architectural implementation of a requirements-oriented computing structure. ACM SIGMICRO Newsletter, 10(4):93–100, December 1979. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014188.803017.

Huck:1979:DES

[186] Jerry Huck and Charles Neuhauser.

I/O device emulation in The Stanford Emulation Laboratory. ACM SIGMI-CRO Newsletter, 10(4):101–108, December 1979. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014188.
803018.

Hartung:1979:VE

[187] R. L. Hartung and A. Ammerman. Virtual I/O, an experiment. ACM SIG-MICRO Newsletter, 10(4):109–113, December 1979. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014188.
803019.

Cooper:1980:DEI

[188] R. E. M. Cooper. The direct execution of intermediate languages on an Eclipse computer. ACM SIGMICRO Newsletter, 11(1):6–9, March 1980. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218011.1218012.

Skordalakis:1980:DMD

[189] E. Skordalakis. DGCAS as a microprogram development tool. ACM SIGMICRO Newsletter, 11(1):10–16, March 1980. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1218011.
1218013.

Tracz:1980:MCAa

[190] William J. Tracz. Microprogrammable computer architecture design considerations. ACM SIGMICRO Newsletter, 11(1):17–20, March 1980. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1218011.1218014.

Tracz:1980:MCAb

[191] William J. Tracz. Microprogrammable computer architecture design considerations. ACM SIGMICRO Newsletter, 11(2):10–14, June 1980. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1096404.1096405.

Hansen:1980:FCA

[192] Igor Hansen and Jacek Leszczylowski. On fundamentals of computer-aided design of firmware. ACM SIGMI-CRO Newsletter, 11(3–4):3–12, Decem-

                                        ber 1980. URL https://dl.acm.org/
                                        doi/10.1145/1014190.802705.

Poe:1980:HGO

[193] Michael D. Poe. Heuristics for the global optimization of microprograms. ACM SIGMICRO Newsletter, 11(3–4):13–22, December 1980. URL https://dl.acm.

                                        org/doi/10.1145/1014190.802706.

Dromard:1980:AIS

[194] D. Dromard and F. Dromard. Algorithm implementation on specific microprogrammed structure. ACM SIGMI-CRO Newsletter, 11(3–4):23–29, Decem-

                                        ber 1980. URL https://dl.acm.org/
                                        doi/10.1145/1014190.802707.

Srimani:1980:SSM

[195] Pradip K. Srimani and Bhabani P.

Sinha. Some studies on microprogram optimization. ACM SIGMICRO Newsletter, 11(3–4):30–37, December

1980. URL https://dl.acm.org/doi/
10.1145/1014190.802708.

Ulrich:1980:DMS

[196] John Wade Ulrich. The derivation of microcode by symbolic execution. ACM SIGMICRO Newsletter, 11(3–4):38–42, December 1980. URL https://dl.acm.

org/doi/10.1145/1014190.802709.

Mueller:1980:FAS

[197] R. A. Mueller. Formalization and automated synthesis of microprograms. ACM SIGMICRO Newsletter, 11(3–4): 45–53, December 1980. URL https:

//dl.acm.org/doi/10.1145/1014190.
802710.

Firth:1980:RST

[198] Neal R. Firth. The role of software tools in the development of the ECLIPSE(R) MV/8000 microcode. ACM SIGMI-CRO Newsletter, 11(3–4):54–58, Decem-

ber 1980. URL https://dl.acm.org/
doi/10.1145/1014190.802711.

Reilly:1980:IME

[199] Paul Reilly, Elizabeth Shanahan, and Steven Staudaher. An implementation of microdiagnostics on the ECLIPSE(R) MV/8000. ACM SIGMICRO Newsletter, 11(3–4):59–63, December 1980.

URL https://dl.acm.org/doi/10.
1145/1014190.802712.

Fisher:1980:WJM

[200] Joseph A. Fisher. 2n-way jump microinstruction hardware and an effective instruction binding method. ACM SIGMI-CRO Newsletter, 11(3–4):64–75, Decem-

ber 1980. URL https://dl.acm.org/
doi/10.1145/1014190.802713.

Tomita:1980:PEI

[201] Shinji Tomita, Kiyoshi Shibayama, Toshiaki Kitamura, and Hiroshi Hagiwara. Performance evaluation and improvement of a dynamically microprogrammable computer with lowlevel parallelism. ACM SIGMICRO Newsletter, 11(3–4):79–89, December

1980. URL https://dl.acm.org/doi/
10.1145/1014190.802714.

Blau:1980:MAE

[202] Jonathan S. Blau, Charles J. Holland, and David L. Keating. The microarchitecture of the ECLIPSE(R) MV/ 8000: Conception and implementation. ACM SIGMICRO Newsletter, 11(3–4): 90–97, December 1980. URL https:

//dl.acm.org/doi/10.1145/1014190.
802715.

Epstein:1980:EMM

[203] David I. Epstein. The ECLIPSE(R) MV/8000 microsequencer. ACM SIG-MICRO Newsletter, 11(3–4):98–105, De-

cember 1980. URL https://dl.acm.
org/doi/10.1145/1014190.802716.

Kraley:1980:DUM

[204] Michael Kraley, Randall Rettberg, Philip Herman, Robert Bressler, and Anthony Lake. Design of a usermicroprogrammable building block. ACM SIGMICRO Newsletter, 11(3–4): 106–114, December 1980. URL https:

//dl.acm.org/doi/10.1145/1014190.
802717.

Guttag:1980:CCR

[205] Karl M. Guttag. Compressing control ROM for VLSI microprogrammed microprocessors. ACM SIGMICRO Newsletter, 11(3–4):115–121, December

1980. URL https://dl.acm.org/doi/
10.1145/1014190.802718.

Cragon:1980:EPS

[206] Harvey G. Cragon. The economics of programmable system components. ACM SIGMICRO Newsletter, 11(3–4): 122–125, December 1980. URL https:

//dl.acm.org/doi/10.1145/1014190.
802719.

Bell:1980:BTT

[207] R. K. Bell, W. D. Bell, T. C. Cooper, and T. K. McFarland. The big three — today’s 16-bit microprocessor. ACM SIGMICRO Newsletter, 11(3–4):126– 138, December 1980. URL https:

//dl.acm.org/doi/10.1145/1014190.
802720.

Sint:1980:SHL

[208] Marleen Sint. A survey of high level microprogramming languages. ACM SIG-MICRO Newsletter, 11(3–4):141–153, December 1980. URL https://dl.acm.

org/doi/10.1145/1014190.802721.

Kurki-Suonio:1980:DAL

[209] R. Kurki-Suonio and J. Heinänen. A data abstraction language based on microprogramming. ACM SIGMICRO Newsletter, 11(3–4):154–161, December

1980. URL https://dl.acm.org/doi/
10.1145/1014190.802722.

Meyers:1980:DML

[210] W. J. Meyers. Design of a microcode link editor. ACM SIGMICRO Newsletter, 11(3–4):165–170, December

1980. URL https://dl.acm.org/doi/
10.1145/1014190.802723.

Baker:1980:HLL

[211] Henry G. Baker and Clinton Parker. High level language programs run ten times faster in microstore. ACM SIGMI-CRO Newsletter, 11(3–4):171–177, De-

cember 1980. URL https://dl.acm.
org/doi/10.1145/1014190.802724.

Berglass:1980:MAH

[212] Gilbert R. Berglass. A meta-assembler for highly-parallel microprogrammable systems. ACM SIGMICRO Newsletter, 11(3–4):181–189, December 1980.

URL https://dl.acm.org/doi/10.
1145/1014190.802725.

Kleinsteiber:1980:IHM

[213] James R. Kleinsteiber. IBM 4341 hardware/microcode trade-off decisions. ACM SIGMICRO Newsletter, 11(3–4): 190–192, December 1980. URL https:

//dl.acm.org/doi/10.1145/1014190.
802726.

Hemphill:1981:RPF

[214] John M. Hemphill. Review of “Principles of Firmware Engineering in Microprogram Control, by Michael Andrews”, Computer Science Press, 1980. ACM SIGMICRO Newsletter, 12(1):23–24, March 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096407.
1096409.

Hemphill:1981:RDS

[215] John M. Hemphill. Review of “Digital System Design with LSI BIT-Slice Logic, by Glenford J. Myers”, John Wiley and Son, 1980. ACM SIGMICRO Newsletter, 12(1):24–25, March 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096407.1096410.

Neuhauser:1981:RFM

[216] Charles Neuhauser. Review of “Firmware, Microprogramming and Restructurable Hardware, by G. Chroust and J. R. Mühlbacher”, North-Holland Publishing Company. ACM SIGMICRO Newsletter, 12(1):26–28, March 1981. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096407.1096411.

Tracz:1981:MSS

[217] William J. Tracz. Microcode support software for the Modular Computer System. ACM SIGMICRO Newsletter, 12(2):14–37, June 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096413.1096414.

Wallach:1981:RAP

[218] Walter Wallach. Recent articles on protection of computer related innovations. ACM SIGMICRO Newsletter, 12(3):10–12, September 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218016.1218017.

Wallach:1981:NAE

[219] Walter Wallach. News about Euromicro. ACM SIGMICRO Newsletter, 12(3):13–15, September 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218016.1218018.

Wallach:1981:ORP

[220] Walter Wallach. Other recent papers. ACM SIGMICRO Newsletter, 12(3):16, September 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1218016.1218019.

Papachristou:1981:HMS

[221] Christos A. Papachristou. Hardware microcontrol schemes using PLAs. ACM SIGMICRO Newsletter, 12(4):3–16, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014192.
                                        802431.

Mueller:1981:CTV

[222] Robert A. Mueller and Gearold R. Johnson. Contrasting translation, verification and synthesis in software and firmware engineering. ACM SIGMI-CRO Newsletter, 12(4):17–22, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014192.
                                        802432.

Kershaw:1981:TIF

[223] John Kershaw. Two implementations of the ‘FLEX’ machine. ACM SIGMI-CRO Newsletter, 12(4):25–37, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014192.
                                        802433.

Habib:1981:UMA

[224] Stanley Habib and Xue-Liang Yang. The use of a meta-assembler to design an M code interpreter on AMD2900 chips. ACM SIGMICRO Newsletter, 12(4):38–50, December 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802434.

Ishikawa:1981:APV

[225] Chiaki Ishikawa, Ken Sakamura, and Mamoru Maekawa. Adaptation and personalization of VLSI-based computer architecture. ACM SIGMICRO Newsletter, 12(4):51–61, December 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802435.

Roskos:1981:TUS

[226] J. Eric Roskos and Robert I. Winner. Toward user sharing of the microprogramming level under UNIX on the Perkin-Elmer 3220. ACM SIGMI-CRO Newsletter, 12(4):67–73, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014192.
802436.

Geyer:1981:DTU

[227] Steven Geyer and Anthony Lake. Development tools for user microprogramming. ACM SIGMICRO Newsletter, 12(4):74–77, December 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802437.

Greenberg:1981:MMA

[228] Kenneth F. Greenberg. The Micro8 microcode assembler. ACM SIGMI-CRO Newsletter, 12(4):78–82, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014192.
802438.

Hansen:1981:MMD

[229] Igor Hansen. MIDAS — microprogram description and analysis system. ACM SIGMICRO Newsletter, 12(4):87, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014192.802439.

Poe:1981:IDL

[230] Michael D. Poe, Ross Goodell, and Simon Steely. Issues of the design of a low level microprogramming language for global microcode compaction. ACM SIGMICRO Newsletter, 12(4):88–94, December 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014192.802440.

Sint:1981:MMD

[231] Marleen Sint. MIDL — a microinstruction description language. ACM SIG-MICRO Newsletter, 12(4):95–106, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014192.
                                        802441.

Bartlett:1981:MMD

[232] Joel F. Bartlett. MicroTAL — a machine-dependent, high-level microprogramming language. ACM SIGMI-CRO Newsletter, 12(4):109–114, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014192.
                                        802442.

Marwedel:1981:RMG

[233] Peter Marwedel. A retargetable microcode generation system for a high-level microprogramming language.

ACM SIGMICRO Newsletter, 12(4): 115–123, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802443.

Klassen:1981:QIH

[234] Alynn Klassen and Subrata Dasgupta.

S*(QM-1): an instantiation of the high level microprogramming language scheme S* for the nanodata QM-1 1. ACM SIGMICRO Newsletter, 12 (4):124–130, December 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802444.

Hobson:1981:HLM

[235] Richard F. Hobson, Patrick Hannon, and Jonathan Thornburg. High-level microprogramming with APL syntax. ACM SIGMICRO Newsletter, 12(4): 131–139, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802445.

Martinez-Carballido:1981:GMW

[236] Jorge Francisco Martinez-Carballido and V. Michael Powers. General microprogram width reduction using generator sets. ACM SIGMICRO Newsletter, 12(4):144–153, December 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802446.

Sheraga:1981:AMG

[237] Robert J. Sheraga and John L. Gieser.

Automatic microcode generation for horizontally microprogrammed processors. ACM SIGMICRO Newsletter, 12(4):154–168, December 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802447.

Nicolau:1981:UOM

[238] Alexandru Nicolau and Joseph A. Fisher. Using an oracle to measure potential parallelism in single instruction stream programs. ACM SIGMI-CRO Newsletter, 12(4):171–182, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014192.
802448.

Rau:1981:SST

[239] B. R. Rau and C. D. Glaeser. Some scheduling techniques and an easily schedulable horizontal architecture for high performance scientific computing. ACM SIGMICRO Newsletter, 12(4): 183–198, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014192.802449.

Rideout:1981:CLC

[240] Douglas J. Rideout. Considerations for local compaction of nanocode for the nanodata QM-1. ACM SIGMI-CRO Newsletter, 12(4):205–214, December 1981. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014192.
802450.

Akonteh:1982:CMD

[241] Ayola N. Akonteh. Characterisation of microinstructions as destructible flows in controlled graphs. ACM SIGMICRO Newsletter, 13(1):21–29, March 1982. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096416.1096417.

Wallach:1982:RAM

[242] Walter A. Wallach. Recent articles from “Microprocessing and Microprogramming”. ACM SIGMICRO Newsletter, 13(2):11–13, September 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218028.1218029.

Sievers:1982:RMC

[243] Michael Sievers. Review of “Microprogrammed Control and Reliable Design of Small Computers, by George D. Kraft and Wing N. Toy”, Prentice Hall, 1981. ACM SIGMICRO Newsletter, 13(2):14–15, September 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218028.1218030.

Belgard:1982:RMM

[244] Rich Belgard. Review of “Microprogramming and Microprocessing by Rich Belgard”, North-Holland, 1981. ACM SIGMICRO Newsletter, 13(2): 16–17, September 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218028.1218031.

Agrawala:1982:RPS

[245] Ashok K. Agrawala. Review of “Probability and Statistics with Reliability Queueing and Computer Science Application by K. S. Trivedi”, Prentice-Hall, 1982. ACM SIGMICRO Newsletter, 13(3):13, September 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218021.1218022.

Neuhauser:1982:RIF

[246] Charles Neuhauser. Review of “Implementing Functions: Microprocessors and Firmware by L. Richter, P. Le Beux, G. Choust and G. Noguez”, North-Holland. ACM SIGMICRO Newsletter, 13(3):14–16, September 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218021.1218023.

Tracz:1982:RBSa

[247] William J. Tracz. Review of “Bit-Slice Microprocessor Design by John Mick and James Brick”, McGraw-Hill Book Company, 1980. ACM SIGMICRO Newsletter, 13(3):17, September 1982. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1218021.1218024.

Tracz:1982:RBSb

[248] William J. Tracz. Review of “Bit-Slice Design: Controllers and ALUs by Donnamarie E. White”, Garland STPM Press, 1981. ACM SIGMICRO Newsletter, 13(3):18, September 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218021.1218025.

Johnson:1982:FEM

[249] Gearold R. Johnson, Robert A. Mueller, and Michael Andrews. Firmware engineering and microprogramming at Colorade State University. ACM SIGMI-CRO Newsletter, 13(3):19–23, September 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1218021.
1218026.

Wilkes:1982:KAP

[250] Maurice V. Wilkes. Keynote address — the processor instruction set. ACM SIGMICRO Newsletter, 13(4):3–5, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014194.
800928.

Frieder:1982:AOC

[251] Gideon Frieder. The architecture and operational characteristics of the VMX host machine. ACM SIGMICRO Newsletter, 13(4):9–16, December 1982. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1014194.800929.

Hennessy:1982:MMA

[252] John Hennessy, Norman Jouppi, Steven Przybylski, Christopher Rowen, Thomas Gross, Forest Baskett, and John Gill. MIPS: a microprocessor architecture. ACM SIGMICRO Newsletter, 13(4): 17–22, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800930.

Gieser:1982:MDT

[253] John L. Gieser and Robert J. Sheraga. Microarchitecture description techniques. ACM SIGMICRO Newsletter, 13(4):23–32, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800931.

Beauchamp:1982:UMD

[254] Robert W. Beauchamp and Neal R.

Firth. UDSYS a microcode development system. ACM SIGMICRO Newsletter, 13(4):35–41, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014194.800932.

Guffin:1982:MLD

[255] Ronald M. Guffin. A microprogramming language-directed microarchitecture. ACM SIGMICRO Newsletter, 13(4):42–49, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014194.800933.

Takahashi:1982:MIT

[256] Kazutoshi Takahashi, Etsuo Takahashi, Tatsushige Bito, Toshinori Aoyama, and Akihiko Yamada. MDS: an improved total system for firmware development. ACM SIGMICRO Newsletter, 13(4):50–56, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014194.800934.

Shriver:1982:TVD

[257] Bruce D. Shriver. Through the video display terminal and what Alice found there. ACM SIGMICRO Newsletter, 13(4):61–76, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1014194.800935.

Tredennick:1982:CM

[258] Nick Tredennick. The “cultures” of microprogramming. ACM SIGMI-CRO Newsletter, 13(4):79–83, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1014194.
                                        800936.

Wang:1982:DM

[259] David T. Wang. Defensive microprogramming. ACM SIGMICRO Newsletter, 13(4):84–88, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800937.

Burke:1982:CSV

[260] Gary R. Burke. Control schemes for VLSI microprocessors. ACM SIGMI-CRO Newsletter, 13(4):91–95, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014194.
800938.

Li:1982:VVM

[261] Tientien Li. A VLSI view of microprogrammed system design. ACM SIG-MICRO Newsletter, 13(4):96–104, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014194.
800939.

Papachristou:1982:MAF

[262] Christos A. Papachristou and Satnam Sing Gambhir. A microsequencer architecture with firmware support for modular microprogramming. ACM SIGMICRO Newsletter, 13(4): 105–113, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800940.

Gross:1982:ODB

[263] Thomas R. Gross and John L. Hennessy. Optimizing delayed branches. ACM SIGMICRO Newsletter, 13(4): 114–120, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800941.

Vegdahl:1982:PCC

[264] Steven R. Vegdahl. Phase coupling and constant generation in an optimizing microcode compiler. ACM SIGMI-CRO Newsletter, 13(4):125–133, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014194.
800942.

Mezzalama:1982:MCM

[265] M. Mezzalama, P. Prinetto, and G. Filippi. Microcode compaction via microblock definition. ACM SIGMI-CRO Newsletter, 13(4):134–142, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014194.
800943.

Jacobs:1982:MCT

[266] Dean Jacobs, Jan Prins, Peter Siegel, and Kenneth Wilson. Monte Carlo techniques in code optimization. ACM SIG-MICRO Newsletter, 13(4):143–148, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014194.
800944.

Holtkamp:1982:FMS

[267] B. Holtkamp and H. Kaestner. A firmware monitor to support vertical migration decisions in the UNIX operating system. ACM SIGMICRO Newsletter, 13(4):153–162, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800945.

Olbert:1982:CMI

[268] A. G. Olbert. Crossing the machine interface. ACM SIGMICRO Newsletter, 13(4):163–170, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800946.

Namjoo:1982:DCT

[269] Masood Namjoo. Design of concurrently testable microprogrammed control units. ACM SIGMICRO Newsletter, 13(4):173–180, December 1982. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800947.

Skibbe:1982:PME

[270] Robert E. Skibbe. PACE — a microprogram evaluation system. ACM SIG-MICRO Newsletter, 13(4):181–191, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1014194.
800948.

Bogong:1982:EMA

[271] Su Bogong and Ralph Grishman.

Emulating an MIMD architecture. ACM SIGMICRO Newsletter, 13(4): 197–199, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1014194.800949.

Vegdahl:1983:NPC

[272] Steven R. Vegdahl. A new perspective on the classical microcode compaction problem. ACM SIGMICRO Newsletter, 14(1):11–14, March 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218033.1218034.

Chroust:1983:WS

[273] G. Chroust. Where to start? ACM SIGMICRO Newsletter, 14(1):15–16, March 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1218033.
1218035.

Papachristou:1983:FEM

[274] Christos A. Papachristou. Firmware engineering and microprogramming at the University of Cincinnati. ACM SIGMICRO Newsletter, 14(2):12–13, June 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1218042.
1218043.

Winner:1983:MVU

[275] R. I. Winner. Microprogramming at Vanderbilt University. ACM SIGMI-CRO Newsletter, 14(2):14–17, June 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1218042.
1218044.

Flynn:1983:SEL

[276] Michael J. Flynn. Stanford Emulation Laboratory. ACM SIGMICRO Newsletter, 14(3):10–17, September 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218037.1218038.

Linn:1983:CPU

[277] Joseph L. Linn, Subrata Dasgupta, and Bruce D. Shriver. The codes project at the University of Southwestern Louisiana. ACM SIGMICRO Newsletter, 14(3):18–21, September 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218037.1218039.

Stevens:1983:MDD

[278] C. R. Stevens. Microprogram documentation with design in mind. ACM SIGMICRO Newsletter, 14(3): 22–29, September 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218037.1218040.

Flynn:1983:SDK

[279] M. J. Flynn. Session details: Keynote address. ACM SIGMICRO Newsletter, 14(4), December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/3262833.

Flynn:1983:TBI

[280] Michael J. Flynn. Towards better instruction sets. ACM SIGMICRO Newsletter, 14(4):3–8, December 1983. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096419.1096421.

Linn:1983:SCG

[281] Joseph L. Linn. SRDAG compaction:

a generalization of trace scheduling to increase the use of global context information. ACM SIGMICRO Newsletter, 14(4):11–22, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096423.

Lah:1983:TCM

[282] Jehkwan Lah and Daniel E. Atkins.

Tree compaction of microprograms. ACM SIGMICRO Newsletter, 14(4): 23–33, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096424.

Henry:1983:LGM

[283] Steven G. Henry, Robert A. Mueller, and Michael Andrews. Local and global microcode compaction using reduction operators. ACM SIGMICRO Newsletter, 14(4):34–43, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096425.

Baba:1983:ETL

[284] Takanobu Baba, Katsuhiro Yamazaki, Nobuyuki Hashimoto, Hiroyuki Kanai, Kenzo Okuda, and Kazuhiko Hashimoto. Experimentation with a two-level microprogrammed multiprocessor computer. ACM SIGMICRO Newsletter, 14(4):47–54, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096427.

Duran:1983:DAM

[285] Jordi Duran and Tulin E. Mangir. A design approach for a microprogrammed control unit with built in self test. ACM SIGMICRO Newsletter, 14(4):55–60, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096428.

Aspinwall:1983:MVM

[286] David B. Aspinwall and Yale N. Patt. Modifications to the VAX-11/780 microarchitecture to support IEEE floating point arithmetic. ACM SIGMI-CRO Newsletter, 14(4):61–69, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096419.
1096429.

Hopkins:1983:HDR

[287] William C. Hopkins. HLLDA defies RISC: thoughts on RISCs, CISCs, and HLLDAs. ACM SIGMICRO Newsletter, 14(4):70–74, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096430.

Fedak:1983:CMF

[288] John F. Fedak. A concurrent microprogramming facility. ACM SIGMI-CRO Newsletter, 14(4):77–83, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096419.
1096432.

Prosser:1983:LED

[289] Franklin Prosser and David Winkel. The logic engine development system support for microprogrammed bit-slice development. ACM SIGMICRO Newsletter, 14(4):84–91, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096433.

Eager:1983:MAR

[290] Michael J. Eager. M29: an advanced retargetable microcode assembler. ACM SIGMICRO Newsletter, 14(4):92–100, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096434.

Wagnon:1983:MW

[291] Gylver Wagnon and Dennis J. W.

Maine. An E-Machine workbench. ACM SIGMICRO Newsletter, 14(4): 101–111, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096435.

Albert:1983:MAI

[292] B. Albert and A. Bode. Microprogrammed associative instructions: results and analysis of a case study in vertical migration. ACM SIGMI-CRO Newsletter, 14(4):115–121, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096419.
1096437.

Shin:1983:IML

[293] H. Shin and M. Malek. Identification of microprogrammable loops for problem oriented architecture synthesis. ACM SIGMICRO Newsletter, 14 (4):122–127, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096438.

Winner:1983:TTO

[294] Robert I. Winner and Edward M. Carter. Toward type-oriented dynamic vertical migration. ACM SIG-MICRO Newsletter, 14(4):128–139, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096419.
1096439.

Schaefer:1983:IPU

[295] Mark T. Schaefer and Yale N. Patt. Improving the performance of UCSD Pascal via microprogramming on the PDP-11/60. ACM SIGMICRO Newsletter, 14(4):140–148, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096440.

Wagner:1983:APR

[296] Alan Wagner and Subrata Dasgupta.

Axiomatic proof rules for a machinespecific microprogramming language. ACM SIGMICRO Newsletter, 14(4): 151–158, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096442.

Mueller:1983:FGM

[297] Robert A. Mueller and Joseph Varghese. Flow graph machine models in microcode synthesis. ACM SIGMI-CRO Newsletter, 14(4):159–167, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096419.
1096443.

Shimizu:1983:MES

[298] Toru Shimizu and Ken Sakamura.

MIXER: an expert system for microprogramming. ACM SIGMICRO Newsletter, 14(4):168–175, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096444.

Gurd:1983:EDM

[299] R. Preston Gurd. Experience developing microcode using a high level language. ACM SIGMICRO Newsletter, 14(4):179–184, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096419.1096446.

Takahashi:1983:AAA

[300] Etsuo Takahashi, Kazutoshi Takahashi, Tatsushige Bito, Tohru Sasaki, and Kazuyoshi Kitano. Automatic address assignment of horizontal microprograms. ACM SIGMICRO Newsletter, 14(4):185–192, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1096419.1096447.

Davidson:1983:HLM

[301] Scott Davidson. High level microprogramming: current usage, future prospects. ACM SIGMICRO Newsletter, 14(4):193–200, December 1983. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1096419.1096448.

Dembinski:1983:ADC

[302] P. Dembinski. On automated design of compacted microprograms. ACM SIG-MICRO Newsletter, 14(4):205–214, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/1096419.
                                        1096450.

Larsen:1983:RRM

[303] Tore Larsen, David Landskov, and Bruce D. Shriver. A resource request model for microcode compaction. ACM SIGMICRO Newsletter, 14(4): 215–226, December 1983. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/1096419.1096451.

Kavi:1984:AQ

[304] Krishna M. Kavi and K. Krishnamohan. Architecture quality. ACM SIGMICRO Newsletter, 15(1):11–19, March 1984. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/1096453.1096454.

Annunzjata:1984:TSM

[305] Massimo Annunzjata, Walter Nosci, and Giacomo R. Sechi. A tool for studying microprogramming. ACM SIGMICRO Newsletter, 15(1):20–25, March 1984. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096453.1096455.

Zambotti:1984:LCV

[306] Larry G. Zambotti and Richard E.

Hardy. LSI-CP: VLSI microprocessor emulates military processors. ACM SIGMICRO Newsletter, 15(1):26–32, March 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096453.
1096456.

Slager:1984:MEM

[307] Jim Slager. Microprocessor employs mainframe performance design techniques. ACM SIGMICRO Newsletter, 15(2):9–14, July 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096458.1096459.

Slager:1984:APT

[308] Jim Slager, Glen Louie, Loys Gindraux, and Bill Rash. Applying pipelining techniques to microprocessors. ACM SIGMICRO Newsletter, 15(2):15–18, July 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096458.1096460.

Slager:1984:IMM

[309] J. Slager, G. Louie, and L. Gindraux.

iAPX 286 microarchitecture to maximize performance. ACM SIGMICRO Newsletter, 15(2):19–21, July 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096458.1096461.

Wilburn:1984:SDS

[310] D. L. Wilburn and John Mick. Step-27 development station. ACM SIGMICRO Newsletter, 15(2):22–36, July 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096458.1096462.

Tracz:1984:RAM

[311] William J. Tracz. Review of “Advances in Microprogramming, by Efrem Mallach and Norman Sondak”, Artech House Inc. 1983. ACM SIGMICRO Newsletter, 15(3):13–15, September 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096464.1096469.

Ris:1984:EAF

[312] Frederic N. Ris. Experience with access functions in an experimental compiler. ACM SIGMICRO Newsletter, 15(3):16–27, September 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096464.1096465.

Hurst:1984:SLP

[313] A. J. Hurst. A source language performance monitoring facility for the B1800 Modula Interpreter. ACM SIGMI-CRO Newsletter, 15(3):28–36, September 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096464.
1096466.

Aeberhard:1984:SCS

[314] John Aeberhard. ‘Soft’ computer speeds new system development. ACM SIGMI-CRO Newsletter, 15(3):37–39, September 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/1096464.
1096467.

Milutinovic:1984:MHC

[315] V. Milutinović, D. Roberts, and K. Hwang. Mapping HLL constructs into microcode for improved execution speed. ACM SIGMICRO Newsletter, 15(4):2–11, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808207.

Hall:1984:MMC

[316] Clifford L. Hall. A microcoded multiprocessor crossbar network communications controller. ACM SIGMICRO Newsletter, 15(4):12–20, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808208.

Fenner:1984:MAE

[317] Jack N. Fenner, Jeffery A. Schmidt, Houssam A. Halabi, and Dharma P. Agrawal. MASCO: an academic exercise in computer design using microprogramming. ACM SIGMICRO Newsletter, 15(4):21–30, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808209.

Wilkes:1984:AVM

[318] J. L. Wilkes. Architecture of a VLSI multiple ISA emulator. ACM SIGMI-CRO Newsletter, 15(4):31–36, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/384281.
808210.

Proulx:1984:APF

[319] David M. Proulx. Applications of pipelining to firmware. ACM SIGMI-CRO Newsletter, 15(4):37–46, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/384281.
808211.

Brown:1984:CSM

[320] John F. Brown and Richard L. Sites. A chip set microarchitecture for a high-performance VAX implementation. ACM SIGMICRO Newsletter, 15(4): 48–54, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808212.

Gries:1984:STU

[321] Robert Gries and James A. Woodward. Software tools used in the development of a VLSI VAX microcomputer. ACM SIGMICRO Newsletter, 15(4):55–58, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808213.

Samudrala:1984:DVV

[322] Sridhar Samudrala, Charles Lo, John F. Brown, and Richard E. Calcagni. Design verification of a VLSI VAX microcomputer. ACM SIGMICRO Newsletter, 15(4):59–63, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808214.

Sherwood:1984:PET

[323] Will Sherwood. A prototype engineering tester for microcode and hardware debugging. ACM SIGMICRO Newsletter, 15(4):64–69, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808215.

Calcagni:1984:PCS

[324] Richard E. Calcagni and Will Sherwood.

Patchable control store for reduced microcode risk in a VLSI VAX microcomputer. ACM SIGMICRO Newsletter, 15(4):70–76, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808216.

Su:1984:ITS

[325] Bogong Su, Shiyuan Ding, and Lan Jin. An improvement of trace scheduling for global microcode compaction. ACM SIGMICRO Newsletter, 15(4): 78–85, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808217.

Papachriston:1984:MAR

[326] Christos A. Papachriston and James M.

Reuter. Microassembly and area reduction techniques for PLA microcode. ACM SIGMICRO Newsletter, 15(4): 86–94, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808218.

Baba:1984:CTL

[327] Takanobu Baba, Mitsuru Ikeda, Katsuhiro Yamazaki, and Kenzo Okuda.

Compaction of two-level microprograms for a multiprocessor computer. ACM SIGMICRO Newsletter, 15(4): 95–104, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808219.

Atkins:1984:IIF

[328] Richard P. Atkins. Improved instruction formation in the exhaustive local microcode compaction algorithm. ACM SIGMICRO Newsletter, 15(4): 105–111, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808220.

Charlton:1984:GSB

[329] C. C. Charlton, D. Jackson, and P. H. Leng. The generation of simulatorbased systems for microcode development. ACM SIGMICRO Newsletter, 15(4):114–121, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808221.

Staas:1984:THM

[330] Gary Staas. TDL: a hardware/ microcode test language interpreter. ACM SIGMICRO Newsletter, 15(4): 122–128, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808222.

Eldridge:1984:MMP

[331] J. Eldridge. A “metasimulator” for microcoded processors. ACM SIGMI-CRO Newsletter, 15(4):129–137, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/384281.
808223.

Holtkamp:1984:ASM

[332] B. Holtkamp and P. Wagner. An algorithm for selection of migration candidates. ACM SIGMICRO Newsletter, 15(4):140–146, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808224.

Heimonen:1984:MII

[333] Juha-Matti Heimonen and Juha Heinanen. Migration implementation by integrating microprogramming and HLL programming. ACM SIGMICRO Newsletter, 15(4):147–154, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/384281.
808225.

Papachristou:1984:AMS

[334] C. A. Papachristou, V. R. Immaneni, and D. B. Sarma. An automatic migration scheme based on modular microcode and structured firmware sequencing. ACM SIGMICRO Newsletter, 15(4):155–164, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808226.

Carter:1984:TMS

[335] Edward M. Carter and Robert I.

Winner. Transparent microprogramming in support of abstract type oriented dynamic vertical migration. ACM SIGMICRO Newsletter, 15(4): 165–178, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808227.

Ardoin:1984:IAR

[336] C. D. Ardoin, J. L. Linn, and B. W. Reynolds. The implementation of the attributed recursive descent architecture in VAX-11/780 microcode. ACM SIGMICRO Newsletter, 15(4): 179–189, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/384281.808228.

Ponder:1984:API

[337] Carl G. Ponder and Yale N. Patt. Alternative proposals for implementing Prolog concurrently and implications regarding their respective microarchitectures. ACM SIGMICRO Newsletter, 15(4):192–203, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/384281.808229.

Tick:1984:SPM

[338] Evan Tick. Sequential Prolog machine: Image and host architectures. ACM SIGMICRO Newsletter, 15(4): 204–216, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/384281.808230.

Dobry:1984:DDI

[339] T. P. Dobry, Y. N. Patt, and A. M. Despain. Design decisions influencing the microarchitecture for a Prolog machine. ACM SIGMICRO Newsletter, 15(4):217–231, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/384281.808231.

Levy:1984:MVU

[340] Beth Levy. Microcode verification using SDVS — the method and a case study. ACM SIGMICRO Newsletter, 15(4):234–245, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808232.

Marcus:1984:SSV

[341] Leo Marcus, Stephen D. Crocker, and Jaisook R. Landauer. SDVS: a system for verifying microcode correctness. ACM SIGMICRO Newsletter, 15(4):246–255, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808233.

Takahashi:1984:NUM

[342] Kazutoshi Takahashi, Etsuo Takahashi, Tatsushige Bitoh, and Takao Sugimoto. A new universal microprogram converter. ACM SIGMICRO Newsletter, 15(4):264–266, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808234.

Marwedel:1984:RCH

[343] Peter Marwedel. A retargetable compiler for a high-level microprogramming language. ACM SIGMICRO Newsletter, 15(4):267–274, December 1984. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808235.

Mueller:1984:GMF

[344] Robert A. Mueller, Joseph Varghese, and Vicki H. Allan. Global methods in the flow graph approach to retargetable microcode generation.

ACM SIGMICRO Newsletter, 15(4): 275–284, December 1984. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/384281.808236.

Bingcong:1985:FPM

[345] Chen Bingcong, Wei Xiaofan, and Yao Xin. On the formal pescription of microoperations and its impact on automatic microcode generation. ACM SIGMICRO Newsletter, 16(2):14–21, May 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218048.1218049.

Hennessy:1985:SSI

[346] John Hennessy. SLIM: a simulation and implementation language for VLSI microcode. ACM SIGMICRO Newsletter, 16(2):22–36, May 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1218048.1218050.

Damm:1985:DSM

[347] W. Damm. Design and specification of microprogrammed computer architectures. ACM SIGMICRO Newsletter, 16(4):3–10, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/18906.18907.

Karplus:1985:EHM

[348] K. Karplus and A. Nicolau. Efficient hardware for multiway jumps and prefetches. ACM SIGMICRO Newsletter, 16(4):11–18, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/18906.18908.

Tudruj:1985:MAQ

[349] M. S. Tudruj and R. F. Gajda. A microprogrammable architecture with quasi time-transparent structured control. ACM SIGMICRO Newsletter, 16(4):19–29, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/18906.18909.

Burkowski:1985:IDD

[350] F. J. Burkowski. An interactive diagnostic/debugging subsystem for bit-slice processors. ACM SIGMICRO Newsletter, 16(4):35–46, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/18906.18910.

Skibbe:1985:PAE

[351] R. E. Skibbe. A practical approach to the evaluation of microcode systems. ACM SIGMICRO Newsletter, 16(4):47–56, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/18906.18911.

Tracz:1985:AMS

[352] W. J. Tracz. Advances in microcode support software. ACM SIGMI-CRO Newsletter, 16(4):57–60, December 1985. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/18906.18912.

Damm:1985:VMC

[353] W. Damm and G. Dohmen. Verification of microprogrammed computer architectures in the S*-system: a case study. ACM SIGMICRO Newsletter, 16(4):61–73, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/18906.18913.

Fagin:1985:CPM

[354] B. Fagin, Y. N. Patt, V. Srini, and A. Despain. Compiling Prolog into microcode: a case study using the NCR/ 32-000. ACM SIGMICRO Newsletter, 16(4):79–88, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/18906.18914.

Woo:1985:AHU

[355] N. S. Woo. The architecture of the hardware unification unit and an implementation. ACM SIGMICRO Newsletter, 16(4):89–98, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/18906.18915.

Patt:1985:HNM

[356] Y. N. Patt, W. M. Hwu, and M. Shebanow. HPS, a new microarchitecture: rationale and introduction. ACM SIGMICRO Newsletter, 16(4): 103–108, December 1985. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/18906.18916.

Patt:1985:CIR

[357] Y. N. Patt, S. W. Melvin, W. M. Hwu, and M. C. Shebanow. Critical issues regarding HPS, a high performance microarchitecture. ACM SIGMI-CRO Newsletter, 16(4):109–116, December 1985. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

                                        acm.org/doi/10.1145/18906.18917.

Catlin:1985:HAL

[358] G. Catlin and B. Paseman. Hardware acceleration of logic simulation using a data flow microarchitecture. ACM SIGMICRO Newsletter, 16(4): 117–123, December 1985. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/18906.18918.

Vegdahl:1985:DIC

[359] S. R. Vegdahl. The design of an interactive compiler for optimizing microprograms. ACM SIGMICRO Newsletter, 16(4):129–136, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/18906.18919.

Hopkins:1985:TIH

[360] W. C. Hopkins, M. J. Horton, and C. S. Arnold. Target-independent highlevel microprogramming. ACM SIGMI-CRO Newsletter, 16(4):137–144, December 1985. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/18906.18920.

Hwang:1985:MDM

[361] J. P.-C. Hwang, C. A. Papachristou, and D. D. Cornett. Microcode development for microprogrammed processors. ACM SIGMICRO Newsletter, 16(4):145–156, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/18906.18921.

Wilburn:1985:SDT

[362] D. L. Wilburn and S. Schleimer. STEP development tools: METASTEP language system. ACM SIGMICRO Newsletter, 16(4):157–164, December 1985. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

                                        acm.org/doi/10.1145/18906.18922.

Patt:1985:MPI

[363] Y. N. Patt and J. K. Ahlstrom. Microcode and the protection of intellectual effort. ACM SIGMICRO Newsletter, 16(4):167–170, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/18906.18923.

Su:1985:SEG

[364] B. Su and S. Ding. Some experiments in global microcode compaction. ACM SIGMICRO Newsletter, 16(4):175–180, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/18906.18924.

Tracz:1985:JJA

[365] W. J. Tracz and B. Boesch. JAM — just another microsequencer. ACM SIGMI-CRO Newsletter, 16(4):181–188, December 1985. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

                                        acm.org/doi/10.1145/18906.18925.

Tudruj:1985:CCS

[366] M. S. Tudruj. A customized control store design in microprogrammed control units. ACM SIGMICRO Newsletter, 16(4):189–199, December 1985. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/18906.18926.

Mueller:1986:FEM

[367] Bob Mueller. Firmware engineering and microprogramming research at Colorado State University. ACM SIG-MICRO Newsletter, 17(1):10–13, March 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/16356.16357.

Denniston:1986:MMV

[368] Brad Denniston. MICROTEC mcASM, version 1.1 — structured microcode assembler. ACM SIGMICRO Newsletter, 17(1):14–18, March 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/16356.16358.

Koyama:1986:MCP

[369] Kiyomi Koyama. Microprogram control of a Prolog machine. ACM SIG-MICRO Newsletter, 17(1):20–32, March 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/16356.16359.

King:1986:ITS

[370] Willis King. IEEE transactions on software engineering call for papers. ACM SIGMICRO Newsletter, 17(2):4, July 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/378397.
378401.

Nicolau:1986:MRP

[371] Alexandru Nicolau. Microprogramming research projects at Cornell University. ACM SIGMICRO Newsletter, 17(2):7, July 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/378397.378402.

Laczko:1986:BCD

[372] Frank Laczko, Bob Myers, and Richard Nawrocki. 32-bit CPU design with the ’AS888/’AS890. ACM SIGMICRO Newsletter, 17(2):8–13, July 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/378397.378403.

Denniston:1986:CBM

[373] Brad Denniston. A comparison between mcASM and META29R. ACM SIGMICRO Newsletter, 17(2):14–18, July 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/378397.
378408.

Rao:1986:AGC

[374] S. Upendra Rao and A. K. Majundar. An algorithm for global compaction of horizontal microprograms. ACM SIGMICRO Newsletter, 17(2):19– 30, July 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/378397.
378412.

Armbruster:1986:MP

[375] Edward Armbruster. MICRO-19 photos. ACM SIGMICRO Newsletter, 17(3):4–5, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/378741.378745.

Vegdahl:1986:RAW

[376] Steven R. Vegdahl. Report on the 19th annual workshop on microprogramming. ACM SIGMICRO Newsletter, 17(3):6–7, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/378741.378749.

Manasiev:1986:FAP

[377] L. Manasiev, A. Petkov, and K. Boyanov. A formal approach to the problems of microcode compaction caused by transient data resources of the microarchitectures. ACM SIGMICRO Newsletter, 17(3):8–12, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/378741.378753.

Rao:1986:PIE

[378] S. Upendra Rao, G. P. Shattacharjee, and A. K. Majumdar. Parallel implementation of an efficient microcode compaction algorithm. ACM SIGMI-CRO Newsletter, 17(3):13–19, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/378741.
378761.

Olangunju:1986:EPI

[379] Amos O. Olangunju and Elvis Borders. Emulators: prospective instruments for instruction in systems programming. ACM SIGMICRO Newsletter, 17(3):20–24, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/378741.378764.

Stanculescu:1986:TMI

[380] Alec G. Stanculescu. Toward machine independent microcode. ACM SIGMI-CRO Newsletter, 17(3):25–32, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/378741.
378766.

Tredennick:1986:IVM

[381] N. Tredennick. The impact of VLSI on microprogramming. ACM SIGMI-CRO Newsletter, 17(4):2–5, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

                                        acm.org/doi/10.1145/19530.19531.

Rosenberg:1986:MDT

[382] J. Rosenberg and D. A. Abramson. Microcode development tools for a capability-based processor. ACM SIG-MICRO Newsletter, 17(4):8–14, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

                                        acm.org/doi/10.1145/19530.19532.

Annunziata:1986:MMS

[383] M. Annunziata, L. Lisca, and G. R. Sechi. MIDETT: Microprogrammedsystem design technique and tools. ACM SIGMICRO Newsletter, 17(4): 15–22, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/19530.19533.

Aiken:1986:DEH

[384] A. Aiken and A. Nicolau. A development environment for horizontal microcode programs. ACM SIGMICRO Newsletter, 17(4):23–31, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/19530.19534.

Nixon:1986:MDL

[385] J. F. Nixon, S. R. Schach, and R. I. Winner. A microarchitecture description language for retargeting firmware tools. ACM SIGMICRO Newsletter, 17(4):34–43, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/19530.19535.

Balakrishnan:1986:ERM

[386] M. Balakrishnan, P. C. P. Bhatt, and B. B. Madan. An efficient retargetable microcode generator. ACM SIGMI-CRO Newsletter, 17(4):44–53, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/19530.19536.

Kuo:1986:ASM

[387] Y. H. Kuo and L. Y. Kung. Algorithmic state machine implementation with hybrid microprocessing/ microprogramming scheme. ACM SIGMICRO Newsletter, 17(4):54–67, December 1986. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/10.
1145/19530.19537.

Gee:1986:IPV

[388] J. Gee, S. W. Melvin, and Y. N.

Patt. The implementation of Prolog via VAX 8600 microcode. ACM SIGMI-CRO Newsletter, 17(4):68–74, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/19530.19538.

Patt:1986:RTG

[389] Y. N. Patt, S. W. Melvin, W. M.

Hwu, M. C. Shebanow, and C. Chen. Run-time generation of HPS microinstructions from a VAX instruction stream. ACM SIGMICRO Newsletter, 17(4):75–81, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/19530.19539.

Koutsougeras:1986:DFG

[390] C. Koutsougeras, C. A. Papachristou, and R. R. Vemuri. Data flow graph partitioning to reduce communication cost. ACM SIGMICRO Newsletter, 17(4):82–91, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/19530.19540.

Su:1986:UEU

[391] B. Su, S. Ding, and J. Xia. URPR — an extension of URCR for software pipelining. ACM SIGMICRO Newsletter, 17(4):94–103, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/19530.19541.

Mueller:1986:CSS

[392] R. A. Mueller, B. Su, M. R. Duda, and B. L. Plomondon. A case study in signal processing microprogramming using the URPR software pipelining technique. ACM SIGMICRO Newsletter, 17(4):104–115, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/19530.19542.

Tracz:1986:SM

[393] W. J. Tracz and R. Belgard. The sociology of microprogramming. ACM SIGMICRO Newsletter, 17(4):118– 122, December 1986. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

                                        URL https://dl.acm.org/doi/10.
                                        1145/19530.19543.

DuBose:1986:MR

[394] D. K. DuBose, D. K. Fotakis, and D. Tabak. A microcoded RISC. ACM SIGMICRO Newsletter, 17(4): 124–128, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/19530.19544.

Franca:1986:DRM

[395] F. M. G. Franca, N. Q. Vasconcelos, and E. S. T. Fernandes. Design and realization of MLM: a multilingual machine. ACM SIGMICRO Newsletter, 17(4):129–137, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/19530.19545.

Abramson:1986:MCB

[396] D. A. Abramson and J. Rosenberg. The microarchitecture of a capability-based computer. ACM SIGMICRO Newsletter, 17(4):138–145, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/19530.19546.

Wahl:1986:DVH

[397] N. J. Wahl, S. R. Schach, and R. I.

Winner. A dynamic very high-level debugger for low-level microprograms. ACM SIGMICRO Newsletter, 17(4): 148–155, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/19530.19547.

Li:1986:BBS

[398] N. Li, Y. Tang, and S. Xu. BBMSS:

a bit-splice based microprogram simulation system. ACM SIGMICRO Newsletter, 17(4):156–161, December 1986. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/19530.19548.

Foster:1986:FBS

[399] J. M. Foster. Formally based static analysis of microcode. ACM SIGMI-CRO Newsletter, 17(4):162–170, December 1986. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

                                        acm.org/doi/10.1145/19530.19549.

Melvin:1986:MBE

[400] S. W. Melvin and Y. N. Patt. A microcode-based environment for noninvasive performance analysis. ACM SIGMICRO Newsletter, 17(4):171– 177, December 1986. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

                                        URL https://dl.acm.org/doi/10.
                                        1145/19530.19550.

Eager:1987:DBS

[401] Michael J. Eager. The death of bitslice. ACM SIGMICRO Newsletter, 18 (1–2):8–10, March 1987. URL https:

                                        //dl.acm.org/doi/10.1145/379531.
                                        379532.

Derwin:1987:ECL

[402] Douglas K. Derwin and Daniel R. Siegel. The effect of copyright law on the development of functionally compatible functionally compatible microprogrammed devices. ACM SIGMI-CRO Newsletter, 18(1–2):11–22, March

                                        1987. URL https://dl.acm.org/doi/
                                        10.1145/379531.379533.

Haverlack:1987:MA

[403] Richard J. Haverlack. Microcode assistant. ACM SIGMICRO Newsletter, 18(3):8–11, September 1987. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/378719.378725.

Burgess:1987:PIS

[404] G. S. Burgess. Preliminary information on SCAMP. ACM SIGMICRO Newsletter, 18(3):12–39, September 1987. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/378719.378734.

Aijazi:1987:CAA

[405] Mahmood G. Aijazi. Computer aided antibiotic therapy. ACM SIGMI-CRO Newsletter, 18(4):13–16, December 1987. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/16360.16361.

Glaser:1987:SEP

[406] John P. Glaser and Jane Edgar. Survey of educational programs in medical information science. ACM SIGMI-CRO Newsletter, 18(4):19–36, December 1987. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/16360.16362.

Allen:1987:PCH

[407] Vivki H. Allen and Robert A. Mueller.

Phase coupling for horizontal microcode generation. ACM SIGMICRO Newsletter, 18(4):24–29, December 1987. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/16360.1096734.

Edwards:1987:HMD

[408] Ken Edwards. HILEVEL microcode development system. ACM SIGMI-CRO Newsletter, 18(4):30–33, December 1987. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/16360.
1096735.

Linn:1987:SME

[409] Joseph L. Linn. A suite of microarchitectures for evaluating microcode compilers for other than for ISA interpretation. ACM SIGMICRO Newsletter, 18(4):34–49, December 1987. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/16360.1096743.

Sayers:1989:HCS

[410] Jerry E. Sayers and David E. Martin. A hypothetical computer to simulate microprogramming and conventional machine language. ACM SIG-MICRO Newsletter, 19–20(4–1):4–10,

                                        March 1989. URL https://dl.acm.
                                        org/doi/10.1145/378818.378829.

Anonymous:1989:JID

[411] Anonymous. JRS: integrated design automation system (IDAS). ACM SIG-MICRO Newsletter, 19–20(4–1):11–17,

                                        March 1989. URL https://dl.acm.
                                        org/doi/10.1145/378818.378833.

Armbruster:1989:MPS

[412] Ed Armbruster. Micro-21 pictorial summary. ACM SIGMICRO Newsletter, 19–20(4–1):18–20, March 1989.

                                        URL https://dl.acm.org/doi/10.
                                        1145/378818.378842.

Anonymous:1989:MTS

[413] Anonymous. Micro-21 technical summary: from the electronic mailbag. ACM SIGMICRO Newsletter, 19–20(4– 1):21–22, March 1989. URL https:

                                        //dl.acm.org/doi/10.1145/378818.
                                        378844.

Patt:1989:MC

[414] Yale Patt. Micro-21 from the Chair. ACM SIGMICRO Newsletter, 19–20 (4–1):23, March 1989. URL https:

                                        //dl.acm.org/doi/10.1145/378818.
                                        378847.

Hwu:1989:MPC

[415] Wen mei Hwu. Micro-21 from the program chair. ACM SIGMICRO Newsletter, 19–20(4–1):24, March 1989.

URL https://dl.acm.org/doi/10.
1145/378818.378848.

Johnson:1989:MI

[416] Gerry Johnson. Micro-22 IRELAND.

ACM SIGMICRO Newsletter, 19–20(4– 1):29–30, March 1989. URL https:

//dl.acm.org/doi/10.1145/378818.
378853.

Wilson:1988:NDP

[417] Fred Wilson. A note on division of positive integers. ACM SIG-MICRO Newsletter, 19(1–2):4, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.1096672.

Okuno:1988:FAF

[418] H. G. Okuno, N. Osato, and I. Takeuchi.

Firmware approach to fast Lisp interpreter. ACM SIGMICRO Newsletter, 19(1–2):5–10, June 1988. URL https:

//dl.acm.org/doi/10.1145/62197.
62198.

Anonymous:1988:DM

[419] Anonymous. Distributed microprogramming. ACM SIGMICRO Newsletter, 19 (1–2):11–12, June 1988. URL https:

//dl.acm.org/doi/10.1145/62197.
62199.

Luque:1988:TAR

[420] E. Luque, T. Sorribes, and A. Ripoll.

Tuning architecture at run-time. ACM SIGMICRO Newsletter, 19(1–2):12–16,

June 1988. URL https://dl.acm.org/
doi/10.1145/62197.62200.

Baba:1988:VMS

[421] T. Baba, H. Minakawa, and K. Okuda. A visual microprogramming system. ACM SIGMICRO Newsletter, 19(1–2): 16–20, June 1988. URL https://dl.

acm.org/doi/10.1145/62197.62201.

Chen:1988:CAD

[422] W. J. Chen and G. N. Reddy. A computer aided design automation system for developing microprogrammed processors: a design approach through HDLs. ACM SIGMI-CRO Newsletter, 19(1–2):20–22, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62202.

Pittman:1988:CAS

[423] R. Pittman and L. Bartel. Computer architecture simulation using a register transfer language. ACM SIGMI-CRO Newsletter, 19(1–2):23–31, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62203.

Bhasker:1988:AMC

[424] Jayaram Bhasker. An algorithm for microcode compaction of VHDL behavioral descriptions. ACM SIGMI-CRO Newsletter, 19(1–2):32–34, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62204.

Su:1988:MCT

[425] B. Su, S. Ding, J. Wang, and J. Xia. Microcode compaction with timing constraints. ACM SIGMI-CRO Newsletter, 19(1–2):34–39, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62205.

Bhasker:1988:CMM

[426] J. Bhasker and T. Samad. Compacting MIMOLA microcode. ACM SIG-MICRO Newsletter, 19(1–2):40–44, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62206.

Howland:1988:TSO

[427] Michael A. Howland, Robert A. Mueller, and Philip H. Sweany. Trace scheduling optimization in a retargetable microcode compiler. ACM SIGMI-CRO Newsletter, 19(1–2):45–49, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62207.

Nowak:1988:GBR

[428] Lothar Nowak. Graph based retargetable microcode compilation in the MIMOLA design system. ACM SIG-MICRO Newsletter, 19(1–2):50–53, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62208.

Uht:1988:CHS

[429] A. K. Uht and J. F. Kolen. On the combination of hardware and software concurrency extraction methods. ACM SIGMICRO Newsletter, 19(1–2):53–57,

June 1988. URL https://dl.acm.org/
doi/10.1145/62197.62209.

Melvin:1988:SMB

[430] S. Melvin and Y. Patt. SPAM: a microcode based tool for tracing operating system events. ACM SIGMI-CRO Newsletter, 19(1–2):58–59, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62210.

Shepherd:1988:UML

[431] David Shepherd. Using mathematical logic and formal methods to write correct microcode. ACM SIGMI-CRO Newsletter, 19(1–2):60–64, June

1988. URL https://dl.acm.org/doi/
10.1145/62197.62211.

Mitchell:1988:IMA

[432] James Mitchell. Implementing a mainframe architecture in a 9370 processor. ACM SIGMICRO Newsletter, 19(3):3–10, September 1988. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/62185.62186.

Kalla:1988:PPU

[433] Ron Kalla. The 9373 and 9375 pipelined processing unit. ACM SIGMI-CRO Newsletter, 19(3):11–15, September 1988. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/62185.62187.

Curley:1988:HDP

[434] Lawrence Curley, James Kuruts, and John Myers. Hardware differences in the 9373 and 9375 processors. ACM SIGMI-CRO Newsletter, 19(3):16–23, September 1988. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/62185.62188.

Maergner:1988:IND

[435] Juergen Maergner and Hartmut R. Schwermer. I370 — a new dimension of microprogramming. ACM SIGMI-CRO Newsletter, 19(3):24–31, September 1988. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/62185.62189.

Su:1988:GMG

[436] Boronr Su, Shiyuan Dink, Jian Vank, and Jinshi Xia. GURPR — a method for global software pipelining. ACM SIGMI-CRO Newsletter, 19(3):32–36, September 1988. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/62185.62190.

Ebcioglu:1988:CTS

[437] Kemal Ebcioğlu. A compilation technique for software pipelining of loops with conditional jumps. ACM SIGMI-CRO Newsletter, 19(3):36–41, September 1988. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/62185.62191.

Walicki:1988:OSR

[438] Jack Walicki and John D. Laughlin.

Operation scheduling in reconfigurable, multifunction pipelines. ACM SIGMI-CRO Newsletter, 19(3):42–45, September 1988. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/62185.62192.

Menzilcioglu:1988:CSU

[439] Onat Menzilcioglu. A case study in using two-level control stores. ACM SIGMI-CRO Newsletter, 19(3):46–48, September 1988. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/62185.62193.

Archer:1988:IPM

[440] David W. Archer. The instruction parsing microarchitecture of CVAX microprocessor. ACM SIGMICRO Newsletter, 19(3):48–51, September 1988. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/62185.62194.

Hwu:1988:EHV

[441] Wen mei W. Hwu and Yale N. Patt. Exploiting horizontal and vertical concurrency via the HPSm microprocessor. ACM SIGMICRO Newsletter, 19(3):52–55, September 1988. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/62185.62195.

Wilson:1988:TMH

[442] James E. Wilson, Steve Melvin, Michael Shebanow, Wen-Mei Hwu, and Yale N. Patt. On tuning the microarchitecture of an HPS implementation of the VAX. ACM SIGMICRO Newsletter, 19(3):56–58, September 1988. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/62185.62196.

Johnson:1989:LML

[443] G. R. Johnson. The landmark microcode legal decision. ACM SIGMI-CRO Newsletter, 20(2):4, June 1989. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/71539.71540.

Gray:1989:J

[444] W. P. Gray. Judgement. ACM SIGMI-CRO Newsletter, 20(2):4–12, June 1989. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/71539.71541.

Milutinovic:1989:BRM

[445] Veljko Milutinovic. Book review: Microprogramming and Firmware Engineering Methods by Stanley Habib, Editor:, Van Nostrand Reinhold, 1988. ACM SIGMI-CRO Newsletter, 20(2):14, June 1989. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/71539.1096726.

Habib:1989:MAS

[446] Stanley Habib. Microprogrammed architectures specified using PAIS-Ley (a description of an ongoing project). ACM SIGMICRO Newsletter, 20(2):15, June 1989. CODEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/71539.1096727.

Wilburn:1989:SES

[447] D. Wilburn. Step engineering: step development system. ACM SIGMICRO Newsletter, 20(2):19–25, June 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/71539.71542.

Gurd:1989:BSS

[448] P. Gurd. Bit slice software: user retargetable microcode tools. ACM SIG-MICRO Newsletter, 20(2):26–31, June 1989. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/71539.71543.

Lall:1989:MIC

[449] K. Lall and J. W. Atwood. A microprogrammed interpreter for Concurrent Euclid. ACM SIGMICRO Newsletter, 20(3):1–10, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75396.

Allan:1989:FLM

[450] S. J. Allan. Functional languages in microcode compilers. ACM SIGMICRO Newsletter, 20(3):11–20, August 1989. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/75395.75397.

Sibai:1989:DPM

[451] F. N. Sibai, L. Watson, and M. Lu. Design and performance measurements of a parallel machine for the unification algorithm. ACM SIGMICRO Newsletter, 20(3):21–30, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/75395.75398.

Brady:1989:DEA

[452] M. Brady. A direct execution architecture for Prolog? ACM SIGMICRO Newsletter, 20(3):31–34, August 1989. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/75395.75399.

Mulder:1989:CED

[453] H. Mulder and R. J. Portier. Costeffective design of application specific VLIW processors using the SCARCE framework. ACM SIGMICRO Newsletter, 20(3):35–42, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/75395.75400.

Nakatani:1989:CCT

[454] T. Nakatani and K. Ebcioğlu. “Combining” as a compilation technique for VLIW architectures. ACM SIGMICRO Newsletter, 20(3):43–55, August 1989. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/75395.75401.

Lenders:1989:MIS

[455] P. M. Lenders, H. Schröder, and P. Strazdins. Microprogramming instruction systolic arrays. ACM SIGMI-CRO Newsletter, 20(3):56–69, August 1989. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https://dl.

acm.org/doi/10.1145/75395.75402.

Parekhji:1989:DMM

[456] R. A. Parekhji and N. K. Nanda. Design methodology and microdiagnostics development for a self-checking microprocessor. ACM SIGMICRO Newsletter, 20(3):70–82, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75403.

Kovacs:1989:EME

[457] L. A. Kovacs and S. F. Gilli. Extended microcode error checking on a pipelined machine. ACM SIGMICRO Newsletter, 20(3):83–87, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75404.

Malaiya:1989:IUU

[458] Y. K. Malaiya. On inherent untestability of unaugmented microprogrammed control. ACM SIGMICRO Newsletter, 20(3):88–96, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75405.

Sechi:1989:ACM

[459] G. R. Sechi. Abstract computing machines. ACM SIGMICRO Newsletter, 20(3):97–111, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75406.

Allan:1989:POT

[460] V. H. Allan. Peephole optimization as a targeting and coupling tool. ACM SIGMICRO Newsletter, 20(3): 112–121, August 1989. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

                                        URL https://dl.acm.org/doi/10.
                                        1145/75395.75407.

Katti:1989:ISL

[461] R. Katti and L. Manwaring. Information structures in language directed architectures. ACM SIGMICRO Newsletter, 20(3):122–126, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/75395.75408.

vandeGoor:1989:DOO

[462] A. J. van de Goor and H. Corporaal. DOAS: an object oriented architecture supporting secure languages. ACM SIGMICRO Newsletter, 20(3): 127–134, August 1989. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

                                        URL https://dl.acm.org/doi/10.
                                        1145/75395.75409.

Liu:1989:LOT

[463] D. Liu and W. K. Giloi. A loop optimization technique based on scheduling table. ACM SIGMICRO Newsletter, 20(3):135–140, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/75395.75410.

Gasperoni:1989:OLP

[464] F. Gasperoni, U. Schwiegelshohn, and K. Ebcioğlu. On optimal loop parallelization. ACM SIGMICRO Newsletter, 20(3):141–147, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75411.

Sanchez:1989:MTE

[465] E. Sanchez. A microprogramming teaching environment using the Macintosh computer. ACM SIGMICRO Newsletter, 20(3):148–155, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75412.

Parker:1989:VBM

[466] A. Parker and J. O. Hamblen. A VLSI based microprogramming evaluation system to support an instructional laboratory. ACM SIGMICRO Newsletter, 20(3):156–159, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75413.

DAngelo:1989:DEA

[467] S. D’Angelo and G. R. Sechi. Definition of elementary arithmetic operations by using ACM. ACM SIGMICRO Newsletter, 20(3):160–162, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75414.

Wijaya:1989:IFL

[468] P. Wijaya and V. H. Allan. Incremental foresighted local compaction. ACM SIGMICRO Newsletter, 20(3): 163–171, August 1989. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/10.
1145/75395.75415.

Linn:1989:AEU

[469] J. L. Linn and C. D. Ardoin. All example of using pseudofields to eliminate version shuffling in horizontal code compaction. ACM SIGMICRO Newsletter, 20(3):172–180, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/75395.75416.

Ertem:1989:MOM

[470] M. C. Ertem. Multiple operation memory structures. ACM SIGMICRO Newsletter, 20(3):181–187, August 1989. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/75395.75417.

Chang:1989:FSC

[471] P.-H. Chang and W.-M. W. Hwu. Forward semantic: a compiler-assisted instruction fetch method for heavily pipelined processors. ACM SIGMICRO Newsletter, 20(3):188–198, August 1989. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/75395.75418.

Shieh:1989:RIS

[472] J.-J. Shieh and C. Papachristou. On reordering instruction streams for pipelined computers. ACM SIGMICRO Newsletter, 20(3):199–206, August 1989. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/75395.75419.

Charlton:1989:FMC

[473] C. C. Charlton, D. Jackson, and P. H. Leng. A functional model of clocked microarchitectures. ACM SIGMICRO Newsletter, 20(3):207–212, August 1989. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/75395.75420.

Patt:1989:MCI

[474] Yale N. Patt. Microarchitecture choices (implementation of the VAX). ACM SIGMICRO Newsletter, 20(3): 213–216, August 1989. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/10.
1145/75395.75421.

Nestor:1989:MMD

[475] J. A. Nestor, B. Soudan, and Z. Mayet.

MIES: a microarchitecture design tool. ACM SIGMICRO Newsletter, 20(3): 217–222, August 1989. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/10.
1145/75395.75422.

Mulder:1989:FVC

[476] H. Mulder and P. Stravers. A flexible VLSI core for an adaptable architecture. ACM SIGMICRO Newsletter, 20(3):223–231, August 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/75395.75423.

Fernandes:1989:MMS

[477] E. S. T. Fernandes. A model for microarchitecture structure evaluation. ACM SIGMICRO Newsletter, 20(3): 232–236, August 1989. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/10.
1145/75395.75424.

Flynn:1989:AM

[478] M. J. Flynn and R. I. Winner. ASIC microprocessors. ACM SIGMICRO Newsletter, 20(3):237–243, August 1989. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/75395.75425.

Bondi:1989:MA

[479] Jim Bondi. Micro-22 awards. ACM SIGMICRO Newsletter, 20(4):3, December 1989. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/378796.
                                        378800.

Anonymous:1989:MSB

[480] Anonymous. Minutes of SIGMICRO Board meeting. ACM SIGMICRO Newsletter, 20(4):4–5, December 1989. CODEN SIGMDJ. ISSN 0163-5751,

                                        1050-916X. URL https://dl.acm.org/
                                        doi/10.1145/378796.378807.

Anonymous:1989:BS

[481] Anonymous. Bylaws of SIGmicro. ACM SIGMICRO Newsletter, 20(4):6–9, December 1989. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

                                        //dl.acm.org/doi/10.1145/378796.
                                        378809.

Anonymous:1989:CPM

[482] Anonymous. Call for papers for Micro-23. ACM SIGMICRO Newsletter, 20(4):10, December 1989. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

                                        916X. URL https://dl.acm.org/doi/
                                        10.1145/378796.378812.

Anonymous:1990:ECC

[483] Anonymous. Epsilon C compiler. ACM SIGMICRO Newsletter, 21(1–2):5–16,

                                        June 1990. URL https://dl.acm.org/
                                        doi/10.1145/1096497.1096498.

Allan:1990:SSP

[484] Vicki H. Allan. Sampling of submitted papers. ACM SIGMICRO Newsletter, 21(3):7–15, September 1990. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096500.1096501.

Nakatani:1991:ULW

[485] Toshio Nakatani and Kemal Ebcioğu.

Using a lookahead window in a compaction-based parallelizing compiler. ACM SIGMICRO Newsletter, 22(1):8–19, January 1991. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/1096503.1096504.

Beaty:1991:MFU

[486] Steven Beaty, Darrell Whitley, and Gearold Johnson. Motivation and framework for using genetic algorithms for microcode compaction. ACM SIGMICRO Newsletter, 22(1):20–27, January 1991. CODEN SIGMDJ. ISSN 0163-5751,

1050-916X. URL https://dl.acm.org/
doi/10.1145/1096503.1096505.

Biswas:1991:BRA

[487] Prasenjit Biswas. Book review: The Art of Computer Systems Performance Analysis — by Raj Jain (ISBN 0471-50336-3, 1991, 685 pages, Price: $52.95 John Wiley & Sons Inc., New York). ACM SIGMICRO Newsletter, 22(2): 3, September 1991. CODEN SIG-MDJ. ISSN 0163-5751, 1050-916X.

URL https://dl.acm.org/doi/10.
1145/127050.1096709.

staff:1991:IA

[488] ACM SIGMICRO Newsletter staff. Issue abstracts. ACM SIGMICRO Newsletter, 22(2):7–16, September 1991. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-

916X. URL https://dl.acm.org/doi/
10.1145/127050.1096710.

Butler:1992:IPV

[489] Michael Butler and Yale Patt. An investigation of the performance of various dynamic scheduling techniques. ACM SIGMICRO Newsletter, 23(1–2):1–9, December 1992. URL https://dl.acm.

                                        org/doi/10.1145/144965.144974.

Theobald:1992:LPP

[490] Kevin B. Theobald, Guang R. Gao, and Laurie J. Hendren. On the limits of program parallelism and its smoothability. ACM SIGMICRO Newsletter, 23(1–2): 10–19, December 1992. URL https:

                                        //dl.acm.org/doi/10.1145/144965.
                                        144977.

Vajapeyam:1992:ILC

[491] Sriram Vajapeyam and Wei-Chung Hsu. On the instruction-level characteristics of scalar code in highly-vectorized scientific applications. ACM SIGMI-CRO Newsletter, 23(1–2):20–28, Decem-

                                        ber 1992. URL https://dl.acm.org/
                                        doi/10.1145/144965.144982.

Chang:1992:EIL

[492] Meng chou Chang, Feipei Lai, and Rung ji Shang. Exploiting instructionlevel parallelism with the conjugate register file scheme. ACM SIGMI-CRO Newsletter, 23(1–2):29–32, Decem-

                                        ber 1992. URL https://dl.acm.org/
                                        doi/10.1145/144965.144986.

Tran:1992:LSM

[493] Thang Tran and Chuan lin Wu. Limitation of superscalar microprocessor performance. ACM SIGMICRO Newsletter, 23(1–2):33–36, December

                                        1992. URL https://dl.acm.org/doi/
                                        10.1145/144965.144991.

Chen:1992:BME

[494] Chien-Ming Chen, Yuun-Yen Chen, and Chung-Ta King. Branch merging for effective exploitation of instructionlevel parallelism. ACM SIGMICRO Newsletter, 23(1–2):37–40, December

1992. URL https://dl.acm.org/doi/
10.1145/144965.144993.

DeGloria:1992:NDS

[495] Alessandro De Gloria, Paolo Faraboschi, and Mauro Olivieri. A nondeterministic scheduler for a software pipelining compiler. ACM SIGMI-CRO Newsletter, 23(1–2):41–44, Decem-

ber 1992. URL https://dl.acm.org/
doi/10.1145/144965.144995.

Mahlke:1992:ECS

[496] Scott A. Mahlke, David C. Lin, William Y. Chen, Richard E. Hank, and Roger A. Bringmann. Effective compiler support for predicated execution using the hyperblock. ACM SIGMI-CRO Newsletter, 23(1–2):45–54, Decem-

ber 1992. URL https://dl.acm.org/
doi/10.1145/144965.144998.

Moon:1992:ERC

[497] Soo-Mook Moon and Kemal Ebcioğlu.

An efficient resource-constrained global scheduling technique for superscalar and VLIW processors. ACM SIGMICRO Newsletter, 23(1–2):55–71, December

1992. URL https://dl.acm.org/doi/
10.1145/144965.145000.

Allen:1992:ERS

[498] V. H. Allen, J. Janardhan, R. M.

Lee, and M. Srinivas. Enhanced region scheduling on a program dependence graph. ACM SIGMICRO Newsletter, 23(1–2):72–80, December

                                        1992. URL https://dl.acm.org/doi/
                                        10.1145/144965.145001.

Wolfe:1992:ECP

[499] Andrew Wolfe and Alex Chanin. Executing compressed programs on an embedded RISC architecture. ACM SIG-MICRO Newsletter, 23(1–2):81–91, De-

                                        cember 1992. URL https://dl.acm.
                                        org/doi/10.1145/144965.145003.

Chen:1992:EAL

[500] William Y. Chen, Roger A. Bringmann, Scott A. Mahlke, Richard E. Hank, and James E. Sicolo. An efficient architecture for loop based data preloading. ACM SIGMICRO Newsletter, 23(1–2):92–101, December

                                        1992. URL https://dl.acm.org/doi/
                                        10.1145/144965.145004.

Fu:1992:SDP

[501] John W. C. Fu, Janak H. Patel, and Bob L. Janssens. Stride directed prefetching in scalar processors. ACM SIGMICRO Newsletter, 23(1–2):102– 110, December 1992. URL https:

                                        //dl.acm.org/doi/10.1145/144965.
                                        145006.

Seznec:1992:CSH

[502] André Seznec and Karl Courtel. Controlling and sequencing a heavily pipelined floating-point operator. ACM SIGMICRO Newsletter, 23(1–2):111– 114, December 1992. URL https:

                                        //dl.acm.org/doi/10.1145/144965.
                                        145008.

Uht:1992:DPI

[503] Augustus K. Uht and Darin B. Johnson. Data path issues in a highly concurrent machine. ACM SIGMICRO Newsletter, 23(1–2):115–118, December

1992. URL https://dl.acm.org/doi/
10.1145/144965.145010.

Su:1992:VAO

[504] Bogong Su, Wei Zhao, Zhizhong Tang, and Stanley Habib. A VLIW architecture for optimal execution of branchintensive loops. ACM SIGMICRO Newsletter, 23(1–2):119–124, December

1992. URL https://dl.acm.org/doi/
10.1145/144965.145011.

Knieser:1992:PCB

[505] Michael J. Knieser and Christos A.

Papachristou. Y-Pipe: a conditional branching scheme without pipeline delays. ACM SIGMICRO Newsletter, 23(1–2):125–128, December 1992.

URL https://dl.acm.org/doi/10.
1145/144965.145015.

Yeh:1992:CIF

[506] Tse-Yu Yeh and Yale N. Patt. A comprehensive instruction fetch mechanism for a processor supporting speculative execution. ACM SIGMICRO Newsletter, 23(1–2):129–139, December

1992. URL https://dl.acm.org/doi/
10.1145/144965.145016.

Beckmann:1992:MSD

[507] Carl J. Beckmann and Constantine D.

Polychronopoulos. Microarchitecture support for dynamic scheduling of acyclic task graphs. ACM SIGMICRO Newsletter, 23(1–2):140–148, December

1992. URL https://dl.acm.org/doi/
10.1145/144965.145791.

Malik:1992:ICA

[508] Nadeem Malik, Richard J. Eickemeyer, and Stamatis Vassiliadis. Interlock collapsing ALU for increased instructionlevel parallelism. ACM SIGMICRO Newsletter, 23(1–2):149–157, December

                                        1992. URL https://dl.acm.org/doi/
                                        10.1145/144965.145794.

Rau:1992:CGS

[509] B. Ramakrishna Rau, Michael S. Schlansker, and P. P. Tirumalai. Code generation schema for modulo scheduled loops. ACM SIGMICRO Newsletter, 23(1–2):158–169, December 1992.

                                        URL https://dl.acm.org/doi/10.
                                        1145/144965.145795.

Warter:1992:EMS

[510] Nancy J. Warter, Grant E. Haab, Krishna Subramanian, and John W. Bockhaus. Enhanced modulo scheduling for loops with conditional branches. ACM SIGMICRO Newsletter, 23(1–2):170– 179, December 1992. URL https:

                                        //dl.acm.org/doi/10.1145/144965.
                                        145796.

Vegdahl:1992:DPT

[511] Stephen R. Vegdahl. A dynamicprogramming technique for compacting loops. ACM SIGMICRO Newsletter, 23(1–2):180–188, December 1992.

                                        URL https://dl.acm.org/doi/10.
                                        1145/144965.145797.

Lenir:1992:EIL

[512] Philip Lenir, R. Govindarajan, and S. S. Nemawarkar. Exploiting instructionlevel parallelism: the multithreaded approach. ACM SIGMICRO Newsletter, 23(1–2):189–192, December 1992.

                                        URL https://dl.acm.org/doi/10.
                                        1145/144965.145798.

Tyson:1992:MMI

[513] Gary Tyson, Matthew Farrens, and Andrew R. Pleszkun. MISC: a Multiple Instruction Stream Computer. ACM SIG-MICRO Newsletter, 23(1–2):193–196, December 1992. URL https://dl.acm.

org/doi/10.1145/144965.145800.

Kiyohara:1992:CSV

[514] Tokuzo Kiyohara and John C. Gyllenhaal. Code scheduling for VLIW/ superscalar processors with limited register files. ACM SIGMICRO Newsletter, 23(1–2):197–201, December 1992.

URL https://dl.acm.org/doi/10.
1145/144965.145802.

Conte:1992:TPM

[515] Thomas M. Conte. Tradeoffs in processor/memory interfaces for superscalar processors. ACM SIGMICRO Newsletter, 23(1–2):202–205, December

1992. URL https://dl.acm.org/doi/
10.1145/144965.145804.

Bray:1992:THB

[516] Brian K. Bray and Michael J. Flynn.

Translation hint buffers to reduce access time of physically-addressed instruction caches. ACM SIGMICRO Newsletter, 23(1–2):206–209, December

1992. URL https://dl.acm.org/doi/
10.1145/144965.145806.

Farrens:1992:MVH

[517] Matthew Farrens, Arvin Park, and Gary Tyson. Modifying VM hardware to reduce address pin requirements. ACM SIGMICRO Newsletter, 23(1–2):210– 213, December 1992. URL https:

//dl.acm.org/doi/10.1145/144965.
145808.

Wilken:1992:TZC

[518] Kent D. Wilken and David W. Goodwin. Toward zero-cost branches using instruction registers. ACM SIGMICRO Newsletter, 23(1–2):214–217, December

                                        1992. URL https://dl.acm.org/doi/
                                        10.1145/144965.145810.